### v1

# The Automorphism Equivariant Hitchin Index

Abstract
Let $\mathbb{T}$ be the one-dimensional complex torus. We consider the action of an automorphism $f$ of a Riemann surface $X$ on the cohomology of the $\mathbb{T}$-equivariant determinant line bundle $\mathcal{L}$ over the moduli space $\mathcal{M}$ of rank two Higgs bundles on $X$ with fixed determinant of odd degree. We define and study the automorphism equivariant Hitchin index $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L},f)$. We prove a formula for it in terms of cohomological pairings of canonical $\mathbb{T}$-equivariant classes of certain moduli spaces of parabolic Higgs bundles over the quotient Riemann surface $X/\langle f\rangle$.
1 Introduction
Let $X$ be a compact Riemann surface of genus $g\geq 2$ with an automorphism $f\in\operatorname{Aut}(X)$ of odd prime order $p$. Let $\Lambda\rightarrow X$ be an $\langle f\rangle$-equivariant holomorphic line bundle of odd degree. Denote by $\mathcal{M}$ the moduli space of stable rank two Higgs bundles on $X$ with fixed determinant $\Lambda$, as introduced in [40, 57] The space $\mathcal{M}$ is a smooth quasi-projective $\mathbb{T}$-variety of dimension $6(g-1)$. Further, $\mathcal{M}$ supports a $\mathbb{T}$-equivariant universal Higgs bundle $(\operatorname{\mathbbm{E}},\mathbb{\Phi})\rightarrow\mathcal{M}\times X.$ We construct in Section 2.1 a canonical $\mathbb{T}\times\langle f\rangle$-equivariant structure on $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$. Assume that $X^{f}\not=\emptyset$ and fix $x_{0}\in X^{f}$. Denote by $\pi_{\mathcal{M}}:\mathcal{M}\times X\rightarrow\mathcal{M}$ the projection. Consider the $\mathbb{T}$-eqivariant determinant line bundle $\mathcal{L}\rightarrow\mathcal{M}$ given by
$\mathcal{L}=\det(\operatorname{\mathbbm{E}}_{\mid\mathcal{M}\times\{x_{0}\}})^{1-g}\otimes\det((\pi_{\mathcal{M}})_{!}(\operatorname{\mathbbm{E}})[1]),$ (1)
where $(\pi_{\mathcal{M}})_{!}(\operatorname{\mathbbm{E}})[1]$ is the $1$-shift of the derived pushforward of $\operatorname{\mathbbm{E}}$. The $\mathbb{T}\times\langle f\rangle$-eqivariant structure on $\operatorname{\mathbbm{E}}$ induces one on $\mathcal{L}^{k}=\mathcal{L}^{\otimes k}$, for all $k\in\mathbb{Z}$. Let $t$ be the standard character of $\mathbb{T}$. For $k,q,m,\in\mathbb{Z},q\geq 0$ denote by $H_{m}^{q}(\mathcal{M},\mathcal{L}^{k})$ the $t^{m}$-weight sub-space of $H^{q}(\mathcal{M},\mathcal{L}^{k})$. These are finite-dimensional $\langle f\rangle$-modules.
Definition 1.1.
The automorphism equivariant Hitchin index is the series
$\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)(t)=\sum_{m\leq 0}\sum_{q\geq 0}(-1)^{q}\operatorname{tr}(f:H_{m}^{q}(\mathcal{M},\mathcal{L}^{k})\rightarrow H_{m}^{q}(\mathcal{M},\mathcal{L}^{k}))\cdot t^{m}.$ (2)
Let $(h,I,J,K)$ be the complete hyper-Kähler metric on $\mathcal{M}$ introduced in [40], for which $I$ is compatible with the algebraic structure. Then $\mathcal{L}$ is a quantum bundle for $(\mathcal{M},\omega_{I})$ in the sense of geometric quantization [47, 58]. Our study of (2) is motivated by non-abelian Hodge theory [24, 25, 40], geometric quantization [47, 58], quantum topology [55, 56, 62, 64] and the goal of generalizing the works [3, 7, 8]. This is explained in detail in Section 1.3.
1.1 The Fixed Locus
We use the $\mathbb{T}$-localization result of Wu [68] and the classical Lefshetz fixed point formula [14] to study (2). In connection hereto, we identify the fixed loci $\mathcal{M}^{f}$ and $\mathcal{M}^{\mathbb{T}\times\langle f\rangle}.$ Denote by $\mathcal{N}\subset\mathcal{M}$ the $\langle f\rangle$-subvariety given by the moduli space of stable holomorphic vector bundles with fixed determinant $\Lambda.$ The fixed locus $\mathcal{N}^{f}$ was identified in [5], and we generalize the method from that article.
For each $x\in X^{f}$, define the $p$’th root of unity $\alpha_{x}\in\mu_{p}$ such that the $f$-action on $\Lambda_{x}$ is given by $\alpha_{x}\cdot\operatorname{Id}$. Denote by $\mathcal{I}_{0}$ the set of maps $i$ from $X^{f}$ to the set of unordered pairs of $p$’th roots of unity, such that $m(i(x))=\alpha_{x}$ for all $x\in X^{f}$, where $m$ is the multiplication map. For all $x\in X^{f}$ let $\sqrt{\alpha}_{x}\in\mu_{p}$ be the square-root of $\alpha_{x}$. We can and will identify the set $\mathcal{I}_{0}$ with the set of all tuples $(i_{x})_{x\in X^{f}}$ of non-negative integers with $i_{x}\leq(p-1)/2$ for all $x\in X^{f}$. Under this identification, the map $i$ corresponding to a tuple $(i_{x})_{x\in X^{f}}$ is given by $x\mapsto[(\sqrt{\alpha}_{x}\zeta^{i_{x}},\sqrt{\alpha}_{x}\zeta^{-i_{x}})]$, where $\zeta=e^{2\pi i/p}$.
Consider the projection $\pi:X\rightarrow\tilde{X}$, where $\tilde{X}=X/\langle f\rangle$ is the quotient Riemann surface. Write $\operatorname{\tilde{\mathcal{D}}}=\pi(X^{f})$ and note that $\pi:X^{f}\rightarrow\operatorname{\tilde{\mathcal{D}}}$ is a bijection. For all $x\in X^{f}$ we will write $\pi(x)=x$. For each $i\in\mathcal{I}_{0}$, let $\operatorname{\tilde{\mathcal{D}}}_{i}=\{x\in\operatorname{\tilde{\mathcal{D}}}:i_{x}\not=0\}$. Let ~$\tilde{\mathcal{M}}_{i}$ be the moduli space [46] of rank two stable parabolic Higgs bundles on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ with fixed determinant and weight function both introduced in Definition 2.3. This is a smooth hyper-Kähler $\mathbb{T}$-variety. Let $\mathcal{I}\subset\mathcal{I}_{0}$ be the subset of $i\in\mathcal{I}_{0}$ such that ~$\tilde{\mathcal{M}}_{i}$ is non-empty. Identify $\langle f\rangle$ with the group $\mu_{p}$ and consider ~$\tilde{\mathcal{M}}_{i}$ as a $\mu_{p}$-variety with the trivial action.
Theorem 1.1.
The set of components of $\mathcal{M}^{f}$ is in bijection with $\mathcal{I}$. Fix $i\in\mathcal{I}$ and write $\mathcal{M}_{i}$ for the corresponding component.
  • The component $\mathcal{M}_{i}$ is a $\mathbb{T}$-subvariety of $\mathcal{M}$ and is characterized by the fact that the eigenvalues of the $f$-action on $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i}\times X^{f}}$ are given by $i.$ The map from Definition 2.5 is an isomorphism of $\mathbb{T}$-varieties $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}.$
  • The triple introduced in equation (56) is a universal $\mathbb{T}$-equivariant parabolic Higgs bundle $(\tilde{\operatorname{\mathbbm{E}}}_{i},\tilde{\mathbb{F}}_{i},\tilde{\mathbb{\Phi}}_{i})\rightarrow\tilde{\mathcal{M}}_{i}\times\tilde{X}.$ We have a $\mathbb{T}$-equivariant isomorphism
    $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i}\times X^{f}}\rightarrow\tilde{\operatorname{\mathbbm{E}}}_{i\mid\tilde{\mathcal{M}}_{i}\times\operatorname{\tilde{\mathcal{D}}}},$ (3)
    covering $\varXi_{i}\times\pi$.
Let $i\in\mathcal{I}$. We equip the restriction of $\tilde{\operatorname{\mathbbm{E}}}_{i}$ to ~$\tilde{\mathcal{M}}_{i}\times\operatorname{\tilde{\mathcal{D}}}$ with the $\mathbb{T}\times\mu_{p}$-equivariant structure induced from the isomorphism given in equation (3). We now present the $\mu_{p}$-weight subbundle decomposition of the restriction of $\tilde{\mathbbm{U}}_{i}=\operatorname{End}_{0}(\tilde{\operatorname{\mathbbm{E}}}_{i})$ to ~$\tilde{\mathcal{M}}_{i}\times\operatorname{\tilde{\mathcal{D}}}$. Towards that end let $x\in\operatorname{\tilde{\mathcal{D}}}$. Write $\alpha_{x}=\zeta^{l_{x}}$ with $l_{x}\in\{0,...,p-1\}$. For $x\notin\operatorname{\tilde{\mathcal{D}}}_{i}$, the $\mu_{p}$-action on $\tilde{\mathbb{U}}_{i,x}$ is trivial. Assume now that $x\in\operatorname{\tilde{\mathcal{D}}}_{i}$. Let $\hat{\zeta}$ be the standard representation of $\mu_{p}$. Consider the line bundle on ~$\tilde{\mathcal{M}}_{i}$ given by $\operatorname{\mathbbm{L}}_{i,x}=\operatorname{Hom}(\tilde{\mathbb{F}}_{i,x},\tilde{\operatorname{\mathbbm{E}}}_{i,x}/\tilde{\mathbb{F}}_{i,x})$ which was introduced in [18]. We have that
$\tilde{\mathbb{U}}_{i,x}=\bigoplus_{\epsilon=-1,0,1}\operatorname{\mathbbm{L}}^{\epsilon}_{i,x}\cdot\hat{\zeta}^{2\epsilon(i_{x}+l_{x})}.$ (4)
Theorem 1.1 is of independent interest. A similar identification was obtained in the work [27] in the form of a bijection between $\mathcal{M}^{f}$ and a moduli space of parabolic Higgs bundles on $\tilde{X}$. However, they work with a moduli space of parabolic principal Higgs bundles, which does not admit a $\mathbb{T}$-action, and they don’t consider the equivariant universal bundle.
1.2 Localization
For all $i\in\mathcal{I}$ write $\mathcal{J}_{i}=\pi_{0}(\tilde{\mathcal{M}}_{i}^{\mathbb{T}})$ and denote the components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$ by {~$\{\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}\}_{j\in\mathcal{J}_{i}}$. These components are determined in Corollary 2.1. Fix $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$. Let $\mathbb{Z}[\mu_{p}]$ denote the representation ring of $\mu_{p}$. Let $S$ denote the multiplicatively closed subset generated by the elements $1-\hat{\zeta}^{l}$ with $l\in\{1,...,p-1\}$. The $K$-theory ring of $\mathbb{T}\times\mu_{p}$-equivariant vector bundles on ~$\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$ (below equivariant means with respect to $\mathbb{T}\times\mu_{p}$) is denoted by $R_{i,j}$, and we have that
$R_{i,j}\cong K^{0}_{\mathbb{T}}(\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j})\otimes\mathbb{Z}[\mu_{p}]\cong\bigoplus_{l=0}^{p-1}K^{0}_{\mathbb{T}}(\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j})\cdot\hat{\zeta}^{l}.$ (5)
Let $\hat{R}_{i,j}=R_{i,j}[[t^{-1}]].$ Denote by $s$ the equivariant operation of taking the sum of symmetric powers, and denote by $\lambda$ the equivariant operation of taking the alternating sum of exterior powers. For every equivariant bundle $W\rightarrow\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$, denote by $W^{+}$ (resp. $W^{-}$) the strictly positive (resp. negative) $\mathbb{T}$-weight subbundle of $W$, and denote by $W^{\mathbb{T}}$ the $\mathbb{T}$-fixed subbundle. When $\lambda((W^{\mathbb{T}})^{*})^{-1}$ exists as a class in $S^{-1}\cdot R_{i,j}$ (see Section 3 on localization), we define the equivariant class
$\Omega(W)=\det(W^{-})\cdot s(W^{-})\cdot s((W^{+})^{*})\cdot\lambda((W^{\mathbb{T}})^{*})^{-1}\in S^{-1}\cdot\hat{R}_{i,j}.$ (6)
For all $x\in X^{f}$ define the integer $n_{x}\in\{1,...,p-1\}$ by the condition that $df_{x}=\zeta^{n_{x}}\cdot\operatorname{Id}$, and write $\tilde{\mathbb{U}}_{i,j,x}$ for the restriction of $\tilde{\mathbb{U}}_{i,x}$ to ~$\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. Define the equivariant class
$\tilde{V}_{i,j}=\sum_{x\in\operatorname{\tilde{\mathcal{D}}}}\tilde{\mathbb{U}}_{i,j,x}\cdot(1-t\cdot\hat{\zeta}^{-n_{x}})\sum_{l=1}^{p-1}l\hat{\zeta}^{-ln_{x}}.$ (7)
Consider the $\mathbb{T}$-equivariant determinant line bundle ~$\tilde{\mathcal{L}}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$. As argued in the proof of Theorem 1.2, we have that $({\varXi_{i}}^{-1})^{*}(\mathcal{L}_{\mid\mathcal{M}_{i}})\cong\tilde{\mathcal{L}}_{i}^{\otimes p}$, and the $\mathbb{T}$-part of the induced equivariant structure agrees with the standard one on ~$\tilde{\mathcal{L}}_{i}^{\otimes p}.$ Denote by $T_{i,j}\tilde{\mathcal{M}}_{i}$ the restriction of $T\tilde{\mathcal{M}}_{i}$ to ~$\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. Let $(\tilde{V}_{i,j})^{\mu_{p}}$ be the projection of $\tilde{V}_{i,j}$ onto the 00’th summand in (5). Denote by $\Phi_{p}$ the $p$’th cyclotomic polynomial. In the proof of Theorem 1.2 we argue the existence of an equivariant class $\tilde{\chi}_{i,j}\in S^{-1}\cdot\hat{R}_{i,j}$ given by the following formula
$\tilde{\chi}_{i,j}$ $=\lambda(T^{*}\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}})\cdot\Omega(T_{i,j}\tilde{\mathcal{M}}_{i}\cdot\Phi_{p}(\hat{\zeta}))\cdot\left(\frac{\Omega(\tilde{V}_{i,j})}{\Omega((\tilde{V}_{i,j})^{\mu_{p}}\cdot\Phi_{p}(\hat{\zeta}))}\right)^{\frac{1}{p}},$ (8)
with the $p$’th root choosen such that the leading term (with respect to the $t^{-1}$-grading) of the equivariant $\operatorname{rank}(\tilde{\chi}_{i,j})$ is equal to the element $\tilde{\zeta}_{i,j}$ introduced in Definition 3.1. Let $\tilde{\nu}_{i,j}$ be the integer introduced in Definition 3.1. Denote by $\operatorname{ch}$ the equivariant Chern character.
Theorem 1.2.
The automorphism equivariant Hitchin index localizes to
$\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)=\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{J}_{i}}(-1)^{\tilde{\nu}_{i,j}}\int_{\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}}\operatorname{ch}(\tilde{\chi}_{i,j}\cdot\tilde{\mathcal{L}_{i}}^{pk})\cdot\operatorname{Td}\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}.$ (9)
The decomposition of $\tilde{\mathbb{U}}_{i,x}$ given in (4) determines an expression for $\tilde{\chi}_{i,j}$ via canonical classes [13, 18]. In a planned follow up article, we will derive from Theorem 1.2 an explicit formula for the series (2). See Remark 2 below.
For each $\xi=(\xi_{1},\xi_{2})\in\mu_{p}^{2}$ define $d_{\xi}$ to be the number of $x\in X^{f}$ such that $\zeta^{-n_{x}}=\xi_{1}$ and $\alpha_{x}=\xi_{2}$. All of these are determined by the $\langle f\rangle$-equivariant structure on $\Lambda$, and we set $d_{\Lambda}=(d_{\xi})_{\xi\in\mu_{p}^{2}}$
Definition 1.2.
The fixed point data of $[\Lambda]\in K_{\langle f\rangle}^{0}(X)$ is the tuple $D_{\Lambda}=(g,d_{\Lambda})$.
Define the integer $l_{0}\in\{0,...,p-1\}$ by $\alpha_{x_{0}}=\zeta^{l_{0}}$, where $x_{0}$ is the marked fixed point entering in the definition of the determinant line bundle $\mathcal{L}\rightarrow\mathcal{M}$ given in equation (1). Note $\lvert X^{f}\rvert=\sum_{\xi}d_{\xi}$. Observe that the genus of $\tilde{X}$ is determined by $D_{\Lambda}$ via Hurwitz formula. The cohomological pairings in equation (9) are determined by the fixed point data and $l_{0}$. Thus we have the following
Corollary 1.1.
The series $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$ is determined by $(D_{\Lambda},l_{0}).$
Consider the case $X^{f}=\emptyset$. In this case the line bundle $\mathcal{L}$ is defined slightly differently than in equation (1), as explained in Section 3.1. Let $\tilde{g}$ be the genus of $\tilde{X}$ and assume that $\tilde{g}\geq 2$. Denote by $\tilde{\mathcal{M}}$ the moduli space of rank two Higgs bundles on $\tilde{X}$ with fixed determinant $\Lambda/\langle f\rangle$ (this quotient is a well-defined line bundle as explained below). In this case the isomorphism from Theorem 1.1 gives an isomorpshim of $\mathbb{T}$-varieties $\mathcal{M}^{f}\cong\tilde{\mathcal{M}}.$ For all $m\in\mathbb{Z},q\in\mathbb{Z}_{+}$, denote by $H_{m}^{q}(\tilde{\mathcal{M}},\tilde{\mathcal{L}^{k}})$ the $t^{m}$-weight sub-space of $H^{q}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})$ and define
$\chi_{\mathbb{T}}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})(t)=\sum_{m\leq 0}\sum_{q\geq 0}(-1)^{q}\dim(H_{m}^{q}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k}))\cdot t^{m}.$ (10)
The series (10) was studied for the moduli stack of Higgs bundles of any rank by the first author, Gukov and Pei in [6] resulting in a beautiful formula in terms of Lie theory, and it was studied for the rank two case by Halpern-Leistner in [35]. Moreover, the series (10) is a refinement of the celebrated Verlinde polynomial, which has been thoroughly studied [42, 59, 63, 66, 71].
Corollary 1.2.
In the case that the fixed locus $X^{f}$ is empty and $\tilde{g}\geq 2$, we have that
$\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)(t)=\chi_{\mathbb{T}}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})(t^{p}).$ (11)
In Section 1.3 below, we explain how this work is connected to TQFT. We emphasize that the rest of the article, apart from Section 4, can be read independently of Section 1.3. Therefore the reader only interested in the algebro-geometric contents of this paper, including the proof of Theorem 1.2, can skip Section 1.3.
1.3 Higgs Bundles and Quantum Topology
Corollary 1.1 connects this work to quantum topology, as we will now explain. Let $G=\operatorname{SU}(n)$, let $k\in\mathbb{Z}_{+}$, and let $\Lambda_{k}$ denote the set of irreducible $G$-representations of level at most $k$. It was envisioned by Witten [64] that level-$k$ quantum Chern-Simons theory with gauge group $G$ form a three-dimensional TQFT [12], and that the evaluation of Jones polynomial of knots [44, 43] at $\exp(2\pi i/(k+n))$ could be given an intrinsic definition as certain expectation values in quantum Chern-Simons theory. A three-dimensional TQFT $Z_{k}$ was subsequently constructed mathematically using surgery presentations of three-manifolds and algebraic and combinatorial means by Reshetikhin and Turaev [55, 56, 62]. This is known as the Witten-Reshetikhin-Turaev TQFT. Essentially, this is a symmetric monoidal functor from the category of compact oriented three-dimensional cobordisms with oriented framed $\Lambda_{k}$-labelled tangles, to the category of finite dimensional Hilbert spaces. In particular, to a compact oriented two-manifold $\Sigma$ with a (possible empty) finite subset of points labelled by $\Lambda_{k}$, the TQFT assign a finite dimensional Hilbert space $Z_{k}(\Sigma)$, together with a projective unitary representation of the mapping class group
$Z_{k}:\operatorname{MCG}(\Sigma)\rightarrow\operatorname{PU}(Z_{k}(\Sigma)).$ (12)
The WRT-TQFT $Z_{k}$ gives powerful invariants: The projective representations (12) were proven to be asymptotically faithful by the first author in the work [1]. Further, the WRT-TQFT $Z_{k}$ is conjectured to detect the volume of hyperbolic three-manifolds [45].
In [64] Witten argued that the TQTF-representation should be realizable by quantization of moduli spaces of flat $G$-connections. The quantization of moduli spaces construction was achieved independently in [15] and [41] , and this will be reviewed below (in the coprime case). By a large body of work culminating in the work of the first author and Ueno [11] and involving conformal field theory [11, 61], skein theory [19, 20] and the work [48], it is now established, that the projective representations constructed by quantization [15, 41] (and introduced below in equation (14)) are equivalent to the projective representations coming from the WRT-TQFT (12).
We now review the quantization construction. Let $\Sigma$ be compact oriented two-manifold of genus at least two with a marked point $x_{0}$ labelled by the element $\lambda\in\Lambda_{k}$, which corresponds to the conjugacy class $C=[e^{2\pi i/n}\operatorname{Id}]$, under the standard bijection between $\Lambda_{k}$ and certain conjugacy classes of $G$. Let $\mathcal{N}_{\operatorname{Flat}}$ denote the moduli space of flat principal $G$-bundles on $\Sigma^{*}=\Sigma\setminus\{x_{0}\}$ with holonomy around the puncture contained in $C$. The space $\mathcal{N}_{\operatorname{Flat}}$ is a compact smooth symplectic manifold [13, 28], and it is the space of classical solutions in Chern-Simons theory [23, 26] on $\Sigma^{*}\times\mathbb{R}$. Further, it supports a pre-quantum line bundle $\mathcal{L}_{\operatorname{CS}}\rightarrow\mathcal{N}_{\operatorname{Flat}}$ in the sense of [47, 58, 67]. Let $\mathcal{T}$ denote Teichmüller space of $\Sigma,$ pick a complex structure $\sigma\in\mathcal{T}$ and let $X_{\sigma}=(\Sigma,\sigma)$ be the associated Riemann surface. By the Narasimhan-Seshadri theorem [51] $\mathcal{N}_{\operatorname{Flat}}$ is symplectomorphic to the Kähler variety given by the moduli space $\mathcal{N}_{\sigma}$ of stable holomorphic vector bundles on $X_{\sigma}$ of rank $n$ and determinant $\mathcal{O}_{X}(-x_{0}).$ This is a smooth projective variety and $\mathcal{L}_{\operatorname{CS}}\rightarrow\mathcal{N}_{\operatorname{Flat}}$ is isomorphic to Quillen’s [54] determinant line bundle $\mathcal{L}_{\sigma}\rightarrow\mathcal{N}_{\sigma}.$ This line bundle is ample, i.e. all higher cohomology groups vanish. Consider the level-$k$ quantization of $\mathcal{N}_{\operatorname{Flat}}$ with respect to this Kähler polarization
$H_{k,\sigma}=H^{0}(\mathcal{N}_{\sigma},\mathcal{L}_{\sigma}^{k}).$ (13)
The family of all such Kähler-quantizations of $\mathcal{N}_{\operatorname{Flat}}$ form a smooth vector bundle $H_{k}\rightarrow\mathcal{T}$, with fibre at $\sigma\in\mathcal{T}$ given by (13). By the independent works [41] and [15] this bundle supports a projectively flat $\operatorname{MCG}(\Sigma)$-invariant connection, called a Hitchin connection in [2]. Fix $\sigma\in\mathcal{T}$ and write $X=X_{\sigma},\mathcal{N}=\mathcal{N}_{\sigma}$ and $\mathcal{L}=\mathcal{L}_{\sigma}$. The monodromy of the Hitchin connection results in a projective representation
$V_{k}:\operatorname{MCG}(\Sigma)\rightarrow\operatorname{PGL}(H^{0}(\mathcal{N},\mathcal{L}^{k})),$ (14)
and this is the aforementioned projective representation, which is projectively equivalent to the WRT-TQFT-representation (12).
Consider an automorphism $f\in\operatorname{Aut}(X)$. In this case, the action $V_{k}(f)$ is given by the action of $f$ on $H^{0}(\mathcal{N},\mathcal{L}^{k})$. Consider the closed oriented three-manifold given by the mapping torus of the automorphism $M_{f}=\Sigma\times I/\sim.$ Let $K\subset M_{f}$ be the knot traced out by the marked point. Label $K$ by $\lambda$. Consider the Witten-Reshetikhin-Turaev quantum invariant $Z_{k}(M_{f},K)\in\mathbbm{C}$. As detailed in the works [3, 8, 7], this invariant can be computed (up to framing correction) by the algebro-geometric formula
$Z_{k}(M_{f},K,\lambda)=\operatorname{tr}(f:H^{0}(\mathcal{N},\mathcal{L}^{k})\rightarrow H^{0}(\mathcal{N},\mathcal{L}^{k}))=\chi(\mathcal{N},\mathcal{L}^{k},f).$ (15)
In the works [3, 8, 7] the index $\chi(\mathcal{N},\mathcal{L}^{k},f)$ was studied (in more general cases with singular moduli spaces of parabolic vector bundles) by a powerful localization technique valid for projective varieties with singularities, developed in [16, 17]. This was used to prove Witten’s asymptotic expansion conjecture [64] for $M_{f}$. For elements $\varphi\in\operatorname{MCG}(\Sigma)$, whose action on $\mathcal{N}_{\operatorname{Flat}}$ satisfy a genericity condition, this conjecture was proven for the mapping torus $M_{\varphi}$ using the representations (14) in the work [10], which generalized a result from [21].
We now specialize to rank two. The index $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$ is a generalization of the index $\chi(\mathcal{N},\mathcal{L}^{k},f)$. Recall that we have a natural inclusion $\mathcal{N}\subset\mathcal{M}^{\mathbb{T}}$, where $\mathcal{M}$ (resp. $\mathcal{N}$) denotes the moduli space of rank two Higgs bundles (resp. holomorphic vector bundles) with fixed determinant of odd degree. The Narasimhan-Seshadri theorem is generalized by non-abelian Hodge theory [24, 25, 40], which asserts that we have a diffeomorphism
$\mathcal{M}\cong\mathcal{M}_{\operatorname{Flat}},$ (16)
where $\mathcal{M}_{\operatorname{Flat}}$ is the space of classical solutions in complex Chern-Simons theory [22], namely the moduli space of flat principal $\operatorname{SL}(2,\mathbbm{C})$-bundles on $\Sigma^{*}$ with holonomy negative one times the identity around the puncture. In fact $\mathcal{M}$ is a complexification of $\mathcal{N}_{\operatorname{Flat}}$ in the sense of brane quantization [34]. In the physics literature [30], this was used in combination with mirror symmetry [38] to illuminate aspects of the Verlinde polynomial. The diffeomorphism (16) is analytic in complex structure $J$, whereas we consider in this article the complex structure $I$. Quantization of $\mathcal{M}_{\operatorname{Flat}}$ with respect to real polarizations have been considered in the literature [4, 29, 65].
Consider again the mapping torus $M_{f}$. This is Seifert fibered [53] over the quotient surface $\tilde{\Sigma}=\Sigma/\langle f\rangle$. We recall the Seifert invariants of $M_{f}$ in equation (121) below. The knot $K$ traced out by the marked point $x_{0}$ is one of the exceptional fibers. By the Seifert invariants of $(M_{f},K)$ we shall mean the Seifert invariants of $M_{f}$ together with the information of which of the Seifert invariants correspond to $K$. Assume that $\Lambda=\mathcal{O}_{X}(-x_{0})$ is the $\langle f\rangle$-equivariant line bundle associated with the divisor $-x_{0}$. As mentioned above, this is the natural choice of fixed determinant arising from gauge theory via non-abelian Hodge theory [24, 25, 40]. In this case, the Seifert invariants are equivalent to $(D_{\Lambda},l_{0})$. Therefore, Corollary 1.1 trivially implies the following
Corollary 1.3.
The series $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$ is determined by the Seifert invariants of the pair $(M_{f},K)$.
Corollary 1.3 is well-motivated by TQFT axioms [12]. Further, we have that
$\lim_{t\rightarrow\infty}(-1)^{3g-3}\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)(t)=\chi(\mathcal{N},\mathcal{L}^{k+2},f).$ (17)
This equation is explained in detail Section 4. Equation (17) provides a new link between algebraic geometry and quantum topology, which extends in a natural way the link provided by the realization of the WRT-TQFT mapping class group representations via quantization of moduli spaces of holomorphic vector bundles on Riemann surfaces.
Finally, we mention that for the class of plumbed three-manifolds, there is the BPS $q$-series invariant $\hat{Z}(M;q)$ [31, 33], which is an integer power series invariant (depending on a choice of $\operatorname{Spin}^{c}$-structure), which is a refinement of $Z_{k}(M)$ and connected to $\operatorname{SL}(2,\mathbbm{C})$-Chern-Simons via resurgence [9, 32]. At the time of writing, it is not clear if and how this is connected to the automorphism equivariant Hitchin index.
1.4 Organization
The organization of this paper is as follows. In Section 2.1 we construct the $\mathbb{T}\times\langle f\rangle$-equivariant structure on $(\operatorname{\mathbbm{E}},\mathbb{\Phi})\rightarrow\mathcal{M}\times X$ and prove Theorem 1.1. In Section 3 we prove Theorem 1.2 and Corollary 1.1 and Corollary 1.2. Corollary 1.3 is proven in Section 4.
Acknowledgements.
The authors warmly thank T. Hausel for very valuable discussions and assistance concerning Higgs bundles and their properties.
This work is supported by the grant from the Simons foundation, Simons Collaboration on New Structures in Low-Dimensional Topology grant no. 994320, the ERC-SyG project, Recursive and Exact New Quantum Theory (ReNewQuantum) with funding from the European Research Council under the European Union‘s Horizon 2020 research and innovation programme, grant agreement no. 810573 and the Carlsberg Foundation grant no. CF20-0431.
2 The Fixed Locus
Remark 1.
Let $\Sigma$ be a compact Riemann surface with a finite subset $\delta$ of marked points. Our conventions and notation on quasi-parabolic Higgs bundles and parabolic Higgs bundles on $(\Sigma,\delta)$ agree with those in [46], except we only consider full flags and specialize to the rank two case. Notice that the real numbers $\{\alpha_{j}(y)\}_{j=1,2,y\in\delta}$ introduced in Section $1$ in [46] satisfy that $\alpha_{j}(y)=w_{j}(y),j=1,2$ for all marked points $y\in\delta$.
Definition 2.1.
Let $D\subset X^{f}$ be a subset. Given a quasi-parabolic Higgs bundle $(E,F,\Phi)\rightarrow(X,D)$ define the quasi-parabolic Higgs bundle $f^{*}(E,F,\Phi)=(f^{*}(E),f^{*}(F),\tilde{f}^{*}(\Phi))$, where $(f^{*}(E),f^{*}(F))\rightarrow(X,D)$ is the pullback quasi-parabolic vector bundle and
$\tilde{f}^{*}(\Phi)=(\operatorname{Id}_{\operatorname{End}_{0}(f^{*}(E))}\otimes(df)^{*})(f^{*}(\Phi)).$ (18)
Consider the case $D=\emptyset$. The map $(E,\Phi)\mapsto f^{*}(E,\Phi)$ preserves stability and descends to an automorphism of $\mathcal{M}$ of order $p$.
Definition 2.2.
An $\langle f\rangle$-equivariant structure on a quasi-parabolic Higgs bundle $(E,F,\Phi)\rightarrow(X,D)$ is an $\langle f\rangle$-equivariant structure on $E\rightarrow X$, such that the generator $f_{E}$ of the action on $E$ covering the action of $f$ on $X$ induces an isomorphism of quasi-parabolic Higgs bundles $(E,F,\Phi)\rightarrow f^{*}(E,F,\Phi),$ which, by abuse of notation, will be denoted by $f_{E}:(E,F,\Phi)\rightarrow f^{*}(E,F,\Phi)$.
Proposition 2.1.
Let $(E,\Phi)$ be a Higgs bundle with isomorphism class $\operatorname{\mathcal{E}}\in\mathcal{M}^{f}$. There is a unique $\langle f\rangle$-equivariant structure with generator $f_{E}:(E,\Phi)\rightarrow f^{*}(E,\Phi)$ covering the action of $f$ on $X$, such that $\det(E)$ is $\langle f\rangle$-equivariantly isomorphic to $\Lambda.$ Denote by $D_{E}$ the subset of $x\in X^{f}$ such that the eigenvalues of $f_{E}(x)$ are disctint, say $\sqrt{\alpha_{x}}\zeta^{i_{x}}$ and $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$ with $i_{x}\in\{1,...,(p-1)/2\}$. For each $x\in D_{E}$ let $F_{x}$ be the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace of $f_{E}(x)$. Define the set of flags $F=\{F_{x}\}_{x\in D_{E}}$. The isomorphism class of the $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E,F,\Phi,f_{E})$ on $(X,D_{E})$ depends only on $\operatorname{\mathcal{E}}$.
Proof.
We begin with the existence part. By assumption there exists an isomorphism of Higgs bundles $f_{E}:(E,\Phi)\rightarrow(f^{*}(E),\tilde{f}^{*}(\Phi))$ such that
$(f^{c}_{E}\otimes\operatorname{Id}_{K})(\Phi)=\tilde{f}^{*}(\Phi)=(\operatorname{Id}_{\operatorname{End}_{0}(f^{*}(E))}\otimes(df)^{*})(f^{*}(\Phi)),$ (19)
where $f^{c}_{E}:\operatorname{End}_{0}(E)\rightarrow\operatorname{End}_{0}(f^{*}(E))$ is the isomorphism given by conjugation with $f_{E}$ and where $\tilde{f}^{*}(\Phi)$ is defined by equation (18). Notice that equation (19) is invariant under $f_{E}\mapsto\alpha\cdot f_{E}$ for any $\alpha\in\mathbbm{C}^{\times}$. Recall that since $(E,\Phi)$ is stable we have $f_{E}^{p}\in\operatorname{Aut}(E,\Phi)\simeq\mathbbm{C}^{\times}$. Thus, we can normalize to ensure $f_{E}^{p}=\operatorname{Id}$ and because of conjugation invariance, the equation (19) remain true. Thus $f^{*}([E,\Phi])=[(E,\Phi)]$ implies that there exists a lift $f_{E}$ of $f$ to $(E,\Phi)$. This is unique up to multiplying by an element of $\mu_{p}$. Now let $h:\det(E)\rightarrow\Lambda$ be an isomorphims (this exists by assumption). Let $f_{\Lambda}$ be the canonical $\langle f\rangle$-equivariant structure on $\Lambda$. Then $f_{1}:=h\circ\det(f_{E})\circ h^{-1}$ and $f_{\Lambda}$ are two lifts that define $\langle f\rangle$-equivariant stuctures on $\Lambda$. Thus $r=f_{\Lambda}^{-1}\circ f_{1}\in\operatorname{Aut}(\Lambda)\simeq\mathbbm{C}^{\times}$ must satisfy $r^{p}=1$. Since $f_{E}\mapsto\alpha\cdot f_{E}$ has the effect that $r\mapsto\alpha^{2}\cdot r$, we see that there is a unique normalization of $f_{E}$ (still defining an equivariant structure) such that $r=1$, which is equivalent to $h$ being an $\langle f\rangle$-equivariant isomorphism.
For the uniqueness part, notice that if $h^{\prime}:\det(E)\rightarrow\Lambda$ is any other isomorphism, then we must have $h^{\prime}=c\cdot h$ for some $c\in\mathbbm{C}^{\times}$. But then $f_{1}^{\prime}=h^{\prime}\circ f_{E}\circ(h^{\prime})^{-1}=f_{1},$ so $r^{\prime}=f_{\Lambda}^{-1}\circ f_{1}^{\prime}$ also satisfies $r^{\prime}=r$. Thus the normalization of $f_{E}$ for which $h$ is an equivariant isomorphism is the same normalization for which $h^{\prime}$ is. ∎
Recall the notation $(\mathbb{E},\mathbb{\Phi})$ for the $\mathbb{T}$-equivariant universal Higgs bundle. The construction of the $\mathbb{T}$-equivariant structure is presented in [39, §3] and [37, §6.3.2]. The universal bundle is well-defined up to tensoring with a $\mathbb{T}$-equivariant line bundle pulled back from $\mathcal{M}$. Set $\mathbbm{U}=\operatorname{End}_{0}(\mathbb{E})$. We have that
$T\mathcal{M}$ $=R^{1}\pi_{*}(\mathbbm{U}\overset{\operatorname{ad}(\mathbb{\Phi})}{\longrightarrow}\mathbbm{U}\cdot\pi^{*}(K)).$ (20)
We consider ×$\mathcal{M}\times X$ a $\mathbb{T}\times\langle f\rangle$-variety, with the $\langle f\rangle$-action induced by the automorphism $F$ given by $F(\mathcal{E},x)=(f^{*}(\mathcal{E}),f^{-1}(x))$ for all $(\mathcal{E},x)\in\mathcal{M}\times X$. Consider the natural projection onto the second factor $\pi_{X}:\mathcal{M}\times X\rightarrow X$.
Proposition 2.2.
There is a unique $\mathbb{T}\times\langle f\rangle$-equivariant structure on $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ such that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$. Let $f_{\operatorname{\mathbbm{E}}}$ denote the generator of the $\langle f\rangle$-action on $\operatorname{\mathbbm{E}}$ covering the action of $F$ on ×$\mathcal{M}\times X$. Through the isomorphism (20), the generator $f_{\operatorname{\mathbbm{E}}}$ induces the standard $\mathbb{T}\times\langle f\rangle$-equivariant structure on $T\mathcal{M}$.
Proof.
Note that via a formula similar to (18), we can equip $F^{*}(\operatorname{\mathbbm{E}})$ with a pullback universal Higgs field, which we denote by $\tilde{F}^{*}(\mathbb{\Phi})$. It follows from a standard argument that $(F^{*}(\operatorname{\mathbbm{E}}),\tilde{F}^{*}(\mathbb{\Phi}))$ is a universal Higgs bundle. Since the universal Higgs bundle is unique up to tensoring with the pullback of a line bundle on $\mathcal{M}$, and since $\operatorname{Aut}(f)$ act trivially on $\operatorname{Pic}(\mathcal{M})$, we see that there exists a $\mathbb{T}$-equivariant isomorphism of universal Higgs bundles
$F_{\operatorname{\mathbbm{E}}}:(\operatorname{\mathbbm{E}},\mathbb{\Phi})\overset{\sim}{\rightarrow}(F^{*}(\operatorname{\mathbbm{E}}),\tilde{F}^{*}(\mathbb{\Phi})).$ (21)
We will argue that $F_{\operatorname{\mathbbm{E}}}$ can be normalized to a new isomorphism which induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. Clearly $F^{p}=\operatorname{Id}$. It follows that $F_{\operatorname{\mathbbm{E}}}^{p}\in\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$. It is well-known that $\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ is one-dimensional. It follows that $F_{\operatorname{\mathbbm{E}}}^{P}=\lambda\cdot\operatorname{Id}$ for some $\lambda\in\mathbbm{C}\setminus\{0\}$. Thus we can replace $F_{\operatorname{\mathbbm{E}}}$ with $f_{\operatorname{\mathbbm{E}}}:=\lambda^{1/p}F_{\operatorname{\mathbbm{E}}}$ to obtain a lift that induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. Since $\gcd(p,2)=1$, the lift can be fixed by demanding that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$.
The fact that $f_{\operatorname{\mathbbm{E}}}$ induces the differential of the standard $\langle f\rangle$-action on $\mathcal{M}$ is a routine verification, which can be done by writing out the definition of the isomorphism (20). ∎
2.1 The Target Moduli Space of Parabolic Higgs Bundles
Let $[\cdot]_{p}:\mathbb{Z}\rightarrow\{0,...,p-1\}$ denote the remainder modulo $p$. For each $x\in X^{f}$, define $m_{x}\in\{1,...,p-1\}$ by the condition that $m_{x}n_{x}=-1\mod p$, where $n_{x}$ was defined in the introduction by the condition that $df_{x}=\zeta^{n_{x}}\cdot\operatorname{Id}$.
For each $x\in X^{f}$, let $b_{x}\in\{0,...,p-1\}$ be the unique solution to $\zeta^{2b_{x}n_{x}}=\alpha_{x}^{-1}$. Define the $\langle f\rangle$-invariant divisor $B=\sum_{x\in X^{f}}b_{x}$, and let $\mathcal{O}_{B}$ be the $\langle f\rangle$-equivariant line bundle associated with $B$, with the standard $\langle f\rangle$-equivariant structure as presented in Proposition 2.3. Then the eigenvalues of the $f$-action on $\Lambda(2B)$ over $X^{f}$ are all equal to unity. It follows that there exists a unique holomorphic line bundle of odd degree $\tilde{\Lambda}\in\operatorname{Pic}(\tilde{X})$ such that $\Lambda(2B)$ is $\langle f\rangle$-equivariantly isomorphic to $\pi^{*}(\Lambda)$ with the canonical $\langle f\rangle$-equivariant structure on the pullback line bundle. Namely, we can define $\tilde{\Lambda}$ as the quotient $\Lambda(2B)/\langle f\rangle$.
Definition 2.3.
Let $i\in\mathcal{I}_{0}$.
  • Let $D_{i}$ be the divisor associated with the set $\{x\in\operatorname{\tilde{\mathcal{D}}}_{i}:[m_{x}i_{x}]_{p}>(p-1)/2\}.$ Define the weights $w_{i}:\operatorname{\tilde{\mathcal{D}}}_{i}\rightarrow\mathbbm{Q}^{2}$ by $w_{i,1}(x)=0$ and $w_{i,2}(x)=[2m_{x}i_{x}]_{p}/p$.
  • Denote by ~$\tilde{\mathcal{M}}_{i}$ (resp. $\tilde{\mathcal{N}}_{i}$) the moduli space [46] of rank two stable parabolic Higgs bundles (resp. vector bundles) on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i},w_{i})$ with fixed determinant $\tilde{\Lambda}_{i}=\tilde{\Lambda}(D_{i}).$ Define ~$\tilde{\mathcal{M}}=\bigsqcup_{i\in\mathcal{I}}\tilde{\mathcal{M}}_{i}.$
For each $i\in\mathcal{I}$, the moduli space ~$\tilde{\mathcal{M}}_{i}$ is constructed as a hyper-Kähler quotient in [46] and proven to be a quasi-projective $\mathbb{T}$-variety in [69, 70] by means of GIT [50].
2.2 Hecke Transformations and Descent
We recall the following fact about $\langle f\rangle$-equivariant line bundles associated with $\langle f\rangle$-invariant divisors.
Proposition 2.3.
Let $(d_{x})_{x\in X^{f}}$ be a tuple of integers, and let $D$ be the associated $f$-invariant Weil divisor given by $D=\sum_{x\in X^{f}}d_{x}\cdot x$. The associated holomorphic line bundle $\mathcal{O}_{D}\rightarrow X$ admits a canonical $\langle f\rangle$-equivariant structure. Let $f_{D}$ be the generator of the $\langle f\rangle$-action on $\mathcal{O}_{D}$ covering the action of $f$ on $X$. For all $x\in X^{f}$ the generator $f_{D}$ satisfies the following equation $f_{D}(x)=\zeta^{d_{x}n_{x}}\cdot\operatorname{Id}$.
The following is an adaption of Lemma $2.6$ in [5], and defines Hecke transformations for $\langle f\rangle$-equivariant quasi-parabolic Higgs bundles.
Proposition 2.4.
Let $L$ be an $\langle f\rangle$-equivariant line bundle on $X$. Let $(E,F,\Phi)$ be an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle on $X$ such that $\det(E)$ is $\langle f\rangle$-equivariantly isomorphic to $L$. Fix $x\in X^{f}$. Assume that $\Phi$ is holomorphic in a neighbourhood of $x$. Assume that the eigenvalues of $f_{E}(x)$ are $1$ and $\zeta^{b_{x}}$, where $b_{x}\in\{1,...,p-1\}$, and assume $F_{x}$ is the the $\zeta^{b_{x}}$-eigenspace. Consider the $\langle f\rangle$-equivariant line bundle $L^{\prime}=L(-x)$ The following holds. There exists an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E^{\prime},F^{\prime},\Phi^{\prime})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $L^{\prime}$, together with a morphism of $\langle f\rangle$-equivariant quasi-parabolic Higgs bundles $\iota:(E^{\prime},F^{\prime},\Phi^{\prime})\rightarrow(E,F,\Phi)$ with the following properties.
  1. 1.
    The morphism $\iota$ restricts to an isomorphism of $\langle f\rangle$-equivariant quasi-parabolic Higgs bundles over the complement of $x$.
  2. 2.
    The image of $\iota(x)$ is equal to the $1$-eigenspace of $f_{E}(x)$. The kernel of $\iota(x)$ is equal to the $\zeta^{b_{x}-n_{x}}$-eigenspace of $f_{E^{\prime}}(x)$, which is equal to $F^{\prime}_{x}.$ The other eigenvalue of $f_{E^{\prime}}(x)$ is equal to unity.
The following implication holds
$\begin{split}&\operatorname{res}(\Phi^{\prime})_{x}\not=0\implies b_{x}=n_{x}.\end{split}$ (22)
This pair $(E^{\prime},\iota)$ can be described via the following short exact sequence of $\langle f\rangle$-equivariant coherent sheaves, where $T_{x}$ is the sky-scraper sheaf supported at $x$ and associated with the vector space $E_{x}/\ker(f_{E}(x)-\operatorname{Id}_{E}(x))$
[…]
Definition 2.4.
With notation as in Proposition 2.4, the $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E^{\prime},F^{\prime},\Phi^{\prime})$ is called the Hecke transform of $(E,F,\Phi)$ at $x$, and we introduce the notation
$\Psi_{x}(E,F,\Phi):=(E^{\prime},F^{\prime},\Phi^{\prime}),$ (24)
We now give the proof of Proposition 2.4.
Proof.
The following facts are already covered in detail in [5]. The sequence (23) defines an $\langle f\rangle$-equivariant quasi parabolic bundle $(E^{\prime},F^{\prime})$ together with a vector bundle morphism $\iota:E^{\prime}\rightarrow E$, which is an isomorphism of $\langle f\rangle$-equivariant quasi parabolic bundles over $X\setminus\{x\}$. Further, one has an $\langle f\rangle$-equivariant isomorphism $\det(E^{\prime})\simeq\det(E)(-x)$. Thus, it remains to construct the Higgs field $\Phi^{\prime}$ and to prove the assertions about it. The Higgs field $\Phi^{\prime}$ is constructed on $X\setminus\{x\}$ by conjugating $\Phi$ by $\iota$. Thus it is clear that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle above $X\setminus\{x\}$, and it remains only to analyze $\Phi^{\prime}$ near $x$. This is done in two steps. First, we show that $\Phi^{\prime}$ extends meromorphically to $X$ with a potential pole of order one at $x$, and with residue being nilpotent with respect to the flag $F^{\prime}_{x}\subset E^{\prime}_{x}$. It immediately follows that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi parabolic Higgs bundle, because $\Phi^{\prime}$ is fixed by the $\langle f\rangle$-action on a dense subspace. In the second step, we verify the implication (22).
First part.
Let $z$ be a local coordinate for $X$ centered at $x$. Choose a local frame $s=(s_{1},s_{2})$ of $E$ near $x$ such that with respect to this frame we have that $f_{E}(0)=\operatorname{diag}(1,\zeta_{b_{x}})$. Define $s^{\prime}=(s_{1}^{\prime},s_{2}^{\prime})=(s_{1},z\cdot s_{2})$. We now consider $\Phi^{\prime}$ near $x$. Let $(s_{1}^{*},s_{2}^{*})$ be the induced frame of $E^{*}$ near $x$, such that $s_{j}^{*}(s_{i})=\delta_{i,j}$ for $i,j\in\{1,2\}$, where $\delta_{i,j}$ is the Kronecker delta. For $i,j=1,2$ let $\phi_{i,j}\in\mathcal{O}_{X,x}$ be the regular function such that $\phi_{i,j}dz=s_{i}^{*}(\Phi(s_{j}))$. With respect to the frames $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ we see that $\iota$ is given by $\operatorname{diag}(1,z)$. With this notation, we have the following matrix decomposition of $\Phi^{\prime}$ with respect to the ordered frame $(s_{1}^{\prime},s_{2}^{\prime})$
$\iota^{-1}\circ\Phi\circ\iota=\begin{pmatrix}\phi_{1,1}&\phi_{1,2}\cdot z\\ \phi_{2,1}\cdot z^{-1}&\phi_{2,2}\end{pmatrix}dz.$ (27)
It is clear from (27) that $\Phi^{\prime}$ extends meromorphically to $X$ with a potential pole of order one at $x$, the residue of which is nilpotent with respect to $E^{\prime}_{x}\supset F^{\prime}_{x}\supset 0$, since $\mathbbm{C}\cdot s^{\prime}_{2}(x)=F^{\prime}_{x}.$
Second part.
Recall that $s_{2}^{\prime}(x)$ generates the one-dimensional subspace given by $\ker(\iota(x))=F^{\prime}_{x}$. Thus equation (19) implies that
$\operatorname{res}(\Phi^{\prime})_{x}=\phi_{2,1}(x)\cdot(s^{\prime}_{2}\otimes(s^{\prime}_{1})^{*})(x)\in\operatorname{Hom}(E^{\prime}_{x},F^{\prime}_{x}).$ (28)
It follows that if $\operatorname{res}(\Phi^{\prime})_{x}\not=0$, then we must have that $\phi_{2,1}(x)\not=0$. As $(E,\Phi)$ is fixed, we see that equation (19) holds for $\Phi$, and at $x$ this evaluates to
$\begin{pmatrix}&\phi_{1,1}(x)&\zeta^{-b_{x}}\phi_{1,2}(x)\\ &\zeta^{b_{x}}\phi_{2,1}(x)&\phi_{2,2}(x)\end{pmatrix}dz=\zeta^{n_{y}}\begin{pmatrix}&\phi_{1,1}(x)&\phi_{1,2}(x)\\ &\phi_{2,1}(x)&\phi_{2,2}(x)\end{pmatrix}dz.$ (29)
By inspecting the left lower entries of the matrices of the above equation, we see that if $\phi_{2,1}(x)\not=0$, then $b_{x}=n_{x}\mod p$. ∎
For future reference, we briefly analyze the generator $f_{E^{\prime}}$ of the $\langle f\rangle$-action on $E^{\prime}$ near $x$. Keep the notation from the proof of Proposition 2.4. Write $f_{E}=(F_{i,j})_{1\leq i,j\leq 2}$ as a matrix of regular functions $F_{i,j}\in\mathcal{O}_{X,x}$ with respect to the frame $(s_{1},s_{2})$ and the pullback frame $(f^{*}(s_{1}),f^{*}(s_{2}))$. Then we have the following equation for $f_{E^{\prime}}$ with respect to the frames $(s^{\prime}_{1},s_{2}^{\prime})$ and $(f^{*}(s_{1}^{\prime}),f^{*}(s_{2}^{\prime}))$
$\begin{split}f_{E^{\prime}}&=\begin{pmatrix}F_{1,1}&F_{1,2}\cdot z\\ F_{2,1}\cdot z^{-1}\zeta^{-n_{x}}&F_{2,2}\cdot\zeta^{-n_{x}}\end{pmatrix}.\end{split}$ (30)
Here we used $f^{*}(\iota)=\operatorname{diag}(1,z\cdot\zeta^{n_{x}})$ with respect to the pair of ordered frames given by $(f^{*}(s_{1}),f^{*}(s_{2}))$ and $(f^{*}(s^{\prime}_{1}),f^{*}(s^{\prime}_{2})).$ Since $F_{2,1}$ vanishes at $x$, we see from equation (30) that $f_{E^{\prime}}$ extends holomorphically to $X$.
The Hecke transformation $\Psi_{x}$ at $x\in X^{f}$ admits an inverse. This fact follows from an adaption of Lemma $2.8$ in [5]. The following lemma will be useful for analysing stability conditions below.
Lemma 2.1.
Let $(E,F,\Phi)$ and $x\in X^{f}$ be as in Proposition 2.4 and let $(E^{\prime},F^{\prime},\Phi^{\prime})$ denote the Hecke transform. There is a bijection between the holomorphic line subbundles of $E$ and holomorphic line subbundles of $E^{\prime}$, which by abuse of notation is denoted by $\Psi_{x}$, and this has the following properties.
  1. 1.
    It inducess a bijection between $\langle f\rangle$-equivariant holomorphic line subbundles of $E$ and $\langle f\rangle$-equivariant holomorphic line subbundles of $E^{\prime}$,
  2. 2.
    It induces a bijection between $\Phi$-invariant holomorphic line subbundles of $E$ and $\Phi^{\prime}$-invariant holomorphic line subbundles of $E^{\prime}$.
For a line subbundle $H\subset E$ we define $\Psi_{x}(H)$ to be equal to $H(-x)$ if $H_{x}\not=\operatorname{Im}(\iota(x))$, otherwise we define $\Psi_{x}(H)$ to be equal to $H$.
For an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}\rightarrow X$ on which $\langle f\rangle$ act as the identity above $X^{f}$, one can define a quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}/\langle f\rangle\rightarrow\tilde{X}$ by taking the quotient. We say $\operatorname{\mathcal{E}}/\langle f\rangle$ is obtained by descent. This is discussed in detail in the proof of Theorem 1.1 below.
For all $i\in\mathcal{I}_{0}$ define $\mathcal{M}_{i}\subset\mathcal{M}^{f}$ to be the $\mathbb{T}$-subvariety of $\mathcal{M}^{f}$ characterized by the fact that the eigenvalues of the $f$-action on $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i}\times X^{f}}$ are given by $i.$ It will follow from the proof of Theorem 1.2 that $\mathcal{M}_{i}$ is non-empty if and only if ~$\tilde{\mathcal{M}}_{i}$ is non-empty, i.e. if and only if $i\in\mathcal{I}$.
Definition 2.5.
Let $i\in\mathcal{I}_{0}$. Define $\mathcal{D}_{i}=\pi^{-1}(\operatorname{\tilde{\mathcal{D}}}_{i})$ and define the $\langle f\rangle$-invariant divisor $S_{i}$ on $X$ by
$S_{i}=\sum\limits_{x\in\mathcal{D}_{i}}[i_{x}m_{x}]_{p}\cdot x+B.$ (31)
Define the $\mathbb{T}$-equivariant morphism $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$ by
$\varXi_{i}(\operatorname{\mathcal{E}})=\left(\prod\limits_{x\in\mathcal{D}_{i}}\Psi_{x}^{[2m_{x}i_{x}]_{p}}(\operatorname{\mathcal{E}}(S_{i}))\right)/\langle f\rangle.$ (32)
It is easily seen that $\varXi_{i}(\mathcal{M}_{i}\cap\mathcal{N})=\tilde{\mathcal{N}}_{i}$, in accordance with the identification of the $\langle f\rangle$-fixed locus of the moduli space $\mathcal{N}$ of stable bundles given in [5], though we have used a slightly different isomorphism here.
Proof of Theorem 1.1.
For all $i\in\mathcal{I}_{0}$ we define
$\Psi_{i}=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{[2m_{x}i_{x}]_{p}}.$ (33)
The proof is divided into three parts. In the first part, we analyze the effect of $\Psi_{i}$ without regard to stability. In the second part, we analyze descent without regard to stability. Finally, in the third part, we complete the proof by taking stability conditions into account.
First part.
Let $i\in\mathcal{I}_{0}$. We introduce the following notations. Denote by $M_{i}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\Lambda,$ which satisfies the following conditions. For all $x\in X^{f}$, the eigenvalues of $f_{E}(x)$ are $\sqrt{\alpha_{x}}\zeta^{i_{x}}$ and $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$, and $F_{x}\subset E_{x}$ is equal to the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace. Denote by $M_{i}^{\prime}$ the set of isomorphism classes of $\langle f\rangle$-equivariant quasi parabolic rank two Higgs bundles $(E,F,\Phi)\rightarrow(X,\mathcal{D}_{i})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $\pi^{*}(\tilde{\Lambda}_{i})$ and such that $f_{E}$ acts as the identity on the fibres of $E$ above $X^{f}$. We will now show that the map given in equation (33) gives a bijection $\Psi_{i}:M_{i}\rightarrow M_{i}^{\prime}$. First of all, we note that the map $\Psi_{i}$ is well-defined since Hecke transformations at different points $x\in\mathcal{D}_{i}$ commutes. Let $(E_{0},F_{0},\Phi)\in M_{i}$. Define the flags $F_{0}(S_{i})=((F_{0})_{x}\otimes(S_{i})_{x})_{x\in X^{f}}$. Then $(E_{1},F_{1},\Phi):=(E_{0}(S_{i}),F_{0}(S_{i}),\Phi)$ canonically defines a $\langle f\rangle$-equivariant quasi parabolic Higgs bundle as in Proposition 2.4 with eigenvalues $\{1,\zeta^{-2i_{x}}\}_{x\in\mathcal{D}_{i}}$. For all $x\in\mathcal{D}_{i}$ set $b_{i}(x)=[2m_{x}i_{x}]_{p}$. Applying at each $x\in\mathcal{D}_{i}$ a total of $b_{i}(x)$ Hecke modifications to $(E_{1},F_{1},\Phi_{1})$, we obtain a new $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E_{2},F_{2},\Phi_{2})=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{b_{i}(x)}(E_{1},\Phi_{1})$. By construction, the $\langle f\rangle$-action is trivial on $E_{\mid X^{f}}$. This fact follows from Proposition 2.4. We now show $\det(E_{2})\simeq\pi^{*}(\Lambda_{i})$. As each Hecke transform at $x$ tensors the determinant by $(-x)$, we see that the determinant of $E_{2}$ is given by
$\begin{split}&\det(E_{2})=\det(E_{1})\left(-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)=\Lambda\left(2S_{i}-\sum_{x\in\mathcal{D}_{i}}b_{i}(x)\cdot x\right)\\ &=\Lambda\left(2B+\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}])\cdot x\right).\end{split}$ (34)
Recall that $D_{i}=\{x\in\operatorname{\tilde{\mathcal{D}}}_{i},[m_{x}i_{x}]>(p-1)/2\}$. If we have $[m_{x}i_{x}]\leq(p-1)/2$ then $x\notin D_{i}$ and we have $2[m_{x}i_{x}]_{p}=[2m_{x}i_{x}]_{p}$. If not, then $\tilde{x}\in D_{i}$ and one can use that $p$ is odd to show that $[2m_{x}i_{x}]_{p}=2[m_{x}i_{x}]_{p}-p.$ Thus
$\begin{split}&\sum_{x\in\mathcal{D}_{i}}(2[m_{x}i_{x}]_{p}-[2m_{x}i_{x}])\cdot x=\sum_{x\in\mathcal{D}_{i}:[m_{x}i_{x}]>(p-1)/2}p\cdot x=\pi^{*}(D_{i}).\end{split}$ (35)
Combining equation (34) with equation (35), and using $\pi^{*}(\tilde{\Lambda})=\Lambda(2B)$, we obtain the desired isomorphism $\det(E_{2})\simeq\pi^{*}(\tilde{\Lambda}_{i}).$ Thus it only remains to verify that $\Phi_{2}$ is holomorphic on $X\setminus\mathcal{D}_{i}$ with a pole of order at most one at each $x\in\mathcal{D}_{i}$ and with residue nilpotent with respect to the flag. For each tuple $a=(a_{x})_{x\in\mathcal{D}_{i}}$ with $0\leq a_{x}<b_{i}(x)$ define
$\operatorname{\mathcal{E}}_{a}=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{a_{x}}(E_{1},F_{1},\Phi_{1}).$ (36)
It follows from the implication (22), that the Higgs field of $\operatorname{\mathcal{E}}_{a}$ is in fact holomorphic on $X$. Thus, it is only after we apply the last product of Hecke modifications that the resulting Higgs field $\Phi_{2}$ aquires a potential pole divisor, which will be a subset of $\mathcal{D}_{i}$ by construction, and each pole will be of order at most one, with residue nilpotent with respect to the flag (this last fact was shown above and follows from (27)). Thus we have shown that $\Psi_{i}$ defines a bijection $M_{i}\rightarrow M_{i}^{\prime}$.
Second part.
Descent for holomorphic vector bundles is discussed in detail in [5]. We now extend descent to Higgs bundles. Denote by $\tilde{M}_{i}$ the set of isomorphism classes of quasi-parabolic rank two Higgs bundles on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ with fixed determinant $\tilde{\Lambda}_{i}$. We will show that pullback with respect to the quotient map $\pi:X\rightarrow\tilde{X}$ induces a bijection $\pi^{*}:\tilde{M}_{i}\rightarrow M_{i}^{\prime}$ with inverse given by descent $\operatorname{\mathcal{E}}\mapsto\operatorname{\mathcal{E}}/\langle f\rangle$. Define the $\langle f\rangle$-invariant divisor $R_{i}=\mathcal{D}_{i}+(1-p)(X^{f}\setminus\mathcal{D}_{i})$. First we observe that there is a canonical $\langle f\rangle$-equivariant isomorphism
$K_{X}(R_{i})\simeq\pi^{*}(K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$ (37)
Let $\operatorname{\mathcal{E}}^{\prime}=(E^{\prime},F^{\prime},\Phi^{\prime})\in M_{i}^{\prime}$. Clearly the quotient $(E^{\prime}/\langle f\rangle,F^{\prime}/\langle f\rangle)\rightarrow\tilde{X}$ is a well-defined quasi parabolic vector bundle. We will show that
$\Phi^{\prime}\in H^{0}(X,\operatorname{End}_{0}(E^{\prime})\otimes K_{X}(R_{i}))^{\langle f\rangle},$ (38)
and by composing with the canonical isomorphism (37), we can define a holomorphic quotient section
$\Phi^{\prime}/\langle f\rangle\in H^{0}(\tilde{X},\operatorname{End}_{0}(E^{\prime}/f)\otimes K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$
Thus the triple $(E^{\prime}/\langle f\rangle,F^{\prime}/\langle f\rangle,\Phi/\langle f\rangle)$ defines a quasi parabolic Higgs bundle on $\tilde{X}$, which we denote by $\operatorname{\mathcal{E}}^{\prime}/\langle f\rangle\in\tilde{M}_{i}.$ We now show (38). By the first part of the proof, we see that this is equivalent to proving that for all $\operatorname{\mathcal{E}}=(E,F,\Phi)\in M_{i}$, we have that
$\Phi\in H^{0}(X,\operatorname{End}_{0}(E)\otimes K_{X}((1-p)(X^{f}\setminus\mathcal{D}_{i}))).$ (39)
Towards the end of proving (39), let $x\in X^{f}\setminus\mathcal{D}_{i}$. Let $z$ be a coordinate centered at $x$, and let $s=(s_{1},s_{2})$ be a local frame of $E$ near $x$. For $i,j\in\{1,2\}$, let $\phi_{i,j}\in\mathcal{O}_{X,x}$ be the uniquely determined holomorphic functions defined near $x$, such that with respect to the frame $s$, we have that $\Phi=\phi dz$, where $\phi=(\phi_{i,j})_{1\leq i,j\leq 2}.$ We must show that
$\min(\operatorname{ord}_{x}(\phi_{i,j})\mid i,j\in\{1,2\})\geq p-1.$ (40)
As $(E,\Phi)$ is $\langle f\rangle$-invariant, equation (19) holds (with $f_{E}$-denoting the canonical $\langle f\rangle$-generator). Recall that $f_{E}^{c}$ denotes conjugation by $f_{E}$. Consider the left hand side of (19). Let $f^{i,j}_{l,k}\in\mathcal{O}_{X,x}$ be the holomorphic functions defined near $x$ such that
$f^{c}_{E}(\Phi)=(\sum_{l,k}f^{i,j}_{l,k}\cdot\phi_{l,k})_{i,j}dz.$ (41)
We now consider the right hand side of (19). We have that
$\begin{split}&(\operatorname{Id}\otimes(df)^{*})(f^{*}(\phi_{i,j}dz))=f^{*}(\phi_{i,j})(df)^{*}(f^{*}dz)\\ &=f^{*}(\phi_{i,j})\frac{\partial f}{\partial z}dz=\zeta^{n_{x}}f^{*}(\phi_{i,j})dz.\end{split}$ (42)
By (19) we have equality between (41) and (42), and this gives
$\zeta^{n_{x}}\phi_{i,j}\circ f=\sum_{l,k}f^{i,j}_{l,k}\cdot\phi_{l,k}.$ (43)
Evaluating at $z=0$ and using that $f_{l,k}^{i,j}(0)=\delta_{l,k}^{i,j},$ (where the right hand side is Kronecker’s delta function) we obtain that $\phi_{i,j}(0)=0$ Assume inductively that we have shown that $\phi^{(m)}_{l,k}(0)=0$ for all $(l,k)$ and all $m=0,1,2,...,n-1$ with $n\leq p-2$. Differentiating equation (43) $n$ times and evaluating at $z=0$ and using that $\frac{\partial f}{\partial z}(z)=\zeta^{n_{x}}$, we get
$\zeta^{n_{x}}\frac{\partial^{n}\phi_{i,j}\circ f}{\partial z^{n}}(0)$ $=\zeta^{n_{x}(n+1)}\phi^{(n)}_{i,j}(0)=\frac{\partial^{n}}{\partial z^{n}}\sum_{l,k}f^{i,j}_{l,k}\cdot\phi_{l,k}(0)$ (44)
$=\sum_{m=0}^{n}\binom{n}{m}(f^{i,j}_{l,k})^{(n-m)}(0)\phi^{(m)}_{l,k}(0)=\phi^{(n)}_{l,k}(0),$ (45)
where for the last equality, we used our induction hypothesis and that $f_{l,k}^{i,j}(0)=\delta_{l,k}^{i,j}$. In particular $\phi^{(n)}_{l,k}(0)=\zeta^{n_{x}(1+n)}\phi^{(n)}_{l,k}(0).$ As $n+1\leq p-1$ we have $\zeta^{n_{x}(n+1)}\neq 1$ and this shows $\phi^{(n)}_{l,k}(0)=0.$
Third part.
By Proposition 2.1 we have a natural inclusion $\mathcal{M}_{i}\subset M_{i}$. We now prove that the composition given by $(\pi^{*})^{-1}\circ\Psi_{i}:M_{i}\rightarrow\tilde{M}_{i}$ restricts to an isomorphism of $\mathbb{T}$-varieties $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$. Observe that $\varXi_{i}$ is equviariant with respect to the actions of $\mathbb{T}$ on source and target. Therefore it remains to show that
  1. 1.
    We have $\operatorname{Im}(\varXi_{i})=\tilde{\mathcal{M}}_{i}$.
  2. 2.
    The set-theoretic maps $\varXi_{i}$ and $\varXi_{i}^{-1}$ are holomorphic.
We start by establishing $\operatorname{Im}(\varXi_{i})=\tilde{\mathcal{M}}_{i}$. Let $\operatorname{\mathcal{E}}\in M_{i}.$ Because of the first part of the proof, it is enough to show that the underlying Higgs bundle is stable if and only if $\varXi_{i}(\operatorname{\mathcal{E}})$ is parabolically stable with respect to the parabolic weight $w_{i}:\mathcal{D}_{i}\rightarrow\mathbbm{Q}^{2}$. Equivalently, we must show that $\operatorname{\mathcal{E}}^{\prime}=\operatorname{\mathcal{E}}(S_{i})$ is stable if and only if $\varXi_{i}(\operatorname{\mathcal{E}})$ is parabolically stable with respect to the parabolic weight $w_{i}:\operatorname{\tilde{\mathcal{D}}}_{i}\rightarrow\mathbbm{Q}^{2}$. Towards that end, we will first use Lemma 2.1 to show that if $H^{\prime}\subset\operatorname{\mathcal{E}}^{\prime}$ is any subbundle of rank one which is preserved by the Higgs field and the $\langle f\rangle$-action, then we have (with $H=H^{\prime}(-S_{i}))$
$\frac{\mu(\operatorname{\mathcal{E}})-\mu(H)}{p}=\frac{\mu(\operatorname{\mathcal{E}}^{\prime})-\mu(H^{\prime})}{p}=\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))-\mu_{\operatorname{Par}}(\varXi_{i}(H)),$ (46)
where we set
$\varXi_{i}(H)=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{[2m_{x}i_{x}]_{p}}(H(S_{i}))/\langle f\rangle=\prod_{x\in\mathcal{D}_{i}}\Psi_{x}^{[2m_{x}i_{x}]_{p}}(H^{\prime})/\langle f\rangle,$
with notation as in Lemma 2.1, where $\Psi_{x}$ was defined for line subbundles. The first equality in equation (46) is trivial and the second is the analog of Lemma $3.2$ in [5]. We now prove the second equality in (46). As $w_{2}(x)=[2m_{x}i_{x}]_{p}/p$ forall $x\in\mathcal{D}_{i}$, we get from the definition of parabolic slope [46] that
$\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))=\frac{\deg(\varXi_{i}(\operatorname{\mathcal{E}}))+\sum_{x\in\mathcal{D}_{i}}[2m_{x}i_{x}]_{p}/p}{2},$ (47)
$\mu_{\operatorname{Par}}(\varXi_{i}(H))=\deg(\varXi_{i}(H))+\sum_{x\in\mathcal{D}_{i}:H^{\prime}_{x}=F^{\prime}_{x}}\frac{[2m_{x}i_{x}]_{p}}{p}.$ (48)
Since $H^{\prime}$ is $\langle f\rangle$-equivariant, we see that for all $x\in\mathcal{D}_{i}$, we have that if $H^{\prime}_{x}\not=\operatorname{Im}(\iota(x))$ with notation as in Lemma 2.1, then we must have $H^{\prime}_{x}=F^{\prime}_{x}.$ Therefore it follows from Lemma 2.1 that
$\Psi_{x}^{[2m_{x}i_{x}]_{p}}(H^{\prime})=H^{\prime}\left(\sum\limits_{x\in\mathcal{D}_{i}:H^{\prime}_{x}=F^{\prime}_{x}}-[2m_{x}i_{x}]_{p}\cdot x\right),$
and therefore
$\begin{split}&p\deg(\varXi_{i}(H))=\deg(H^{\prime})-\sum_{x\in\mathcal{D}_{i}:H_{x}^{\prime}=F^{\prime}_{x}}[2m_{x}i_{x}]_{p}.\end{split}$ (49)
We have
$p\deg(\varXi_{i}(\operatorname{\mathcal{E}}))=\deg(E^{\prime})-\sum_{x\in\mathcal{D}_{i}}[2m_{x}i_{x}]_{p}.$ (50)
Substituting equation (49) into equation (47) and equation (49) into equation (48) and then computing the difference $\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))-\mu_{\operatorname{Par}}(\varXi_{i}(H))$ gives
$\mu_{\operatorname{Par}}(\varXi_{i}(\operatorname{\mathcal{E}}))-\mu_{\operatorname{Par}}(\varXi_{i}(H))=\frac{\deg(\varXi_{i}(\operatorname{\mathcal{E}}))+\sum_{x\in\mathcal{D}_{i}}[2m_{x}i_{x}]_{p}/p}{2}$ (51)
$-\left(\deg(\varXi_{i}(H))+\sum_{x\in\mathcal{D}_{i}:H^{\prime}_{x}=F^{\prime}_{x}}\frac{[2m_{x}i_{x}]_{p}}{p}\right)$ (52)
$=\frac{\deg(E^{\prime})-\sum_{x\in\mathcal{D}_{i}}[2m_{x}i_{x}]_{p}+\sum_{x\in\mathcal{D}_{i}}[2m_{x}i_{x}]_{p}}{2p}$ (53)
$-\left(\frac{\deg(H^{\prime})-\sum\limits_{x\in\mathcal{D}_{i}:H_{x}^{\prime}=F^{\prime}_{x}}[2m_{x}i_{x}]_{p}}{p}+\sum\limits_{x\in\mathcal{D}_{i}:H^{\prime}_{x}=F^{\prime}_{x}}\frac{[2m_{x}i_{x}]_{p}}{p}\right)$ (54)
$=\frac{\mu(\operatorname{\mathcal{E}})-\mu(H)}{p}.$ (55)
Thus we have proven equation (46). We now use this to prove $\operatorname{Im}(\varXi_{i})=\tilde{\mathcal{M}}_{i}$. Observe that for any $(E,F,\Phi)\in M_{i}^{\prime}$, there is an induced bijection between $\Phi$-invariant and $\langle f\rangle$-equivariant line subbundles $H$ of $E$ and $\Phi/\langle f\rangle$-invariant line subbundles of $E/\langle f\rangle$, and this is given by $H\mapsto H/\langle f\rangle$. This fact and equation (46) implies $\operatorname{Im}(\varXi_{i})\subset\tilde{\mathcal{M}}_{i}.$ Now let $\operatorname{\mathcal{E}}=(E,\Phi)\in M_{i}$ and assume $\varXi_{i}(\operatorname{\mathcal{E}})\in\tilde{\mathcal{M}}_{i}$. Assume towards a contradiction that $\operatorname{\mathcal{E}}$ is not stable. Since $\deg(\Lambda)$ is odd, this is equivalent to $\operatorname{\mathcal{E}}$ not being semi-stable. Consider the Higgs Narasimhan-Harder filtration $0\subsetneq H\subsetneq E$. Observe that $f_{E}(f^{*}(H))$ is a line subbundle which also induce a Higgs Narasimhan-Harder filtration, and therefore we must have $f_{E}(f^{*}(H))=H$ by uniqueness. Therefore $H$ is $\langle f\rangle$-equivariant and $\Phi$-invariant. Hence we can apply equation (46) to obtain a contradiction. Thus $\operatorname{\mathcal{E}}$ is stable. We now argue that $\varXi_{i}$ and its inverse are morphisms of varieties. Both the target and the source of $\varXi_{i}$ can be obtained as GIT quotients [57, 69, 70], and therefore it suffices to show that one can apply Hecke transformations in families. This was done for stable bundles in [5], and the analysis therein can readily be repeated for Higgs bundles. In particular, we can apply Hecke modications to $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ and we define
$(\tilde{\operatorname{\mathbbm{E}}}_{i},\tilde{\mathbb{F}}_{i},\mathbb{\Phi}_{i})=\Psi_{i}((\operatorname{\mathbbm{E}},\mathbb{\Phi})_{\mid\mathcal{M}_{i}\times X})/\langle f\rangle.$ (56)
It is a routine verification that $(\tilde{\operatorname{\mathbbm{E}}}_{i},\tilde{\mathbb{F}}_{i},\mathbb{\Phi}_{i})\rightarrow\tilde{\mathcal{M}}_{i}\times\tilde{X}$ is a $\mathbb{T}$-equivariant universal parabolic Higgs bundle, and that we have a canonical equivariant isomorphism as in equation (3). ∎
2.3 Identification of The Torus Fixed Locus
Definition 2.6.
Let $M$ denote the set of odd integers $m$ with $31\leq m\leq 2g-$ 3. For each $m\in M$, let $\operatorname{\mathcal{F}}_{m}$ be the set of isomorphism classes of rank two Higgs bundles $(E,\Phi_{b})$ on $X$ of the following form.
  • There are holomorphic line bundles $L_{1}$ and $L_{2}$ giving a direct sum decomposition $E=L_{1}\oplus L_{2}$, and such that $L_{1}L_{2}\simeq\Lambda$.
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(X,L_{2}L_{1}^{-1}K_{X})$. The degree of the zero divisor of $b$ is equal to $m$.
For $m\in\mathbb{Z}_{\geq 0}$ let $X_{m}$ (resp. $\tilde{X}_{m}$) denote the space of positive effectice divisors of degree $m$ on $X$ (resp. $\tilde{X}$). Now fix $m\in M$. The following facts are shown by Hitchin in [40]. Every Higgs bundle in $\operatorname{\mathcal{F}}_{m}$ is stable, and thus $\operatorname{\mathcal{F}}_{m}$ is a naturally a subspace of $\mathcal{M}$. Each $\operatorname{\mathcal{F}}_{m}$ is connected and we have
$\mathcal{M}^{\mathbb{T}}=\mathcal{N}\sqcup_{m\in M}\operatorname{\mathcal{F}}_{m}.$ (57)
Moreover, the map $\delta:\operatorname{\mathcal{F}}_{m}\rightarrow X_{m}$, given by $(E,\Phi_{b})\mapsto(b)\in X_{m}$, gives $\operatorname{\mathcal{F}}_{m}$ the structure of a principal $\operatorname{Jac}(X)[2]$-bundle (with action given by tensoring). We notice that pullback with respect to $f$ naturally equips $\operatorname{\mathcal{F}}_{m},\ \operatorname{Jac}(X)[2]$ and $X_{m}$ with an $\langle f\rangle$-action.
Fix $m\in M$. Recall [39] that the universal Higgs bundle restricts to a direct sum of $\mathbb{T}$-equivariant line bundles $\mathbb{E}_{m,1}\oplus\mathbb{E}_{m,0}\rightarrow\operatorname{\mathcal{F}}_{m}\times X$, where $\mathbb{T}$ acts with weight $t$ on $\mathbb{E}_{m,1}$ and with weight equal to unity on $\mathbb{E}_{m,0}$. Fix $i\in\mathcal{I}$. We will now identity the components of $\mathcal{M}_{i}\cap\operatorname{\mathcal{F}}_{m}$ as well as the $\mathbb{T}\times\langle f\rangle$-weight subbundle decomposition of $\operatorname{\mathbbm{E}}$ restricted to these components.
Definition 2.7.
Set $\mathcal{J}_{m}=\sqcup_{i\in\mathcal{I}}\mathcal{J}_{i,m}$, where for each $i\in\mathcal{I}$, we define $\mathcal{J}_{i,m}$ to be the set of divisors of the form $j=\sum_{x\in X^{f}}j_{x}\cdot x,$ such that $l_{j}=(m-\deg(j))/p$ is a non-negative integer and for all $x\in X^{f}$, we have that $j_{x}\in\{0,...,p-1\}$ and
$(2i_{x})^{2}=(n_{x}(1+j_{x}))^{2}\mod p.$ (58)
Let $i\in\mathcal{I},m\in M$ and $j\in\mathcal{J}_{i,m}$. For all $x\in X^{f}$, it follows from equation (58) that there is a unique $\epsilon_{j,x}\in\{1,-1\}$ such that $\epsilon_{j,x}2i_{x}=-n_{x}(1+j_{x})\mod p.$
Lemma 2.2.
The set of components of the $\langle f\rangle$-fixed locus $X_{m}^{f}$ is in bijection with $\mathcal{J}_{m}$, and the set of components of $\mathcal{M}_{i}\cap\mathcal{F}_{m}$ is in bijection with $\mathcal{J}_{i,m}.$ Fix $j\in\mathcal{J}_{i,m}$ and denote the associated components by $X^{f}_{m,j}$ and $\mathcal{M}^{\mathbb{T}}_{i,j}$ respectively.
  1. 1.
    We have an isomorphism $X^{f}_{m,j}\simeq\tilde{X}_{l_{j}}.$ Further, the morphism $\delta$ restricts to a projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle $\mathcal{M}^{\mathbb{T}}_{i,j}\rightarrow X^{f}_{m,j}$. Thus
    $\mathcal{M}_{i,j}^{\mathbb{T}}=\{(E,\Phi_{b})\in\mathcal{M}^{f}\cap\mathcal{F}_{m}:[\operatorname{ord}_{x}(b)]_{p}=j_{x},\ \forall x\in X^{f}\}.$ (59)
  2. 2.
    For all $x\in X^{f}$ we have the equivariant decomposition
    $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i,j}^{\mathbb{T}}\times\{x\}}=\operatorname{\mathbbm{E}}_{m,1}(x)\cdot\sqrt{\hat{\alpha}}_{x}\hat{\zeta}^{\epsilon_{j,x}i_{x}}\oplus\operatorname{\mathbbm{E}}_{m,0}(x)\cdot\sqrt{\hat{\alpha}}_{x}\hat{\zeta}^{-\epsilon_{j,x}i_{x}}.$ (60)
Proof.
$1$ The construction of the isomorphism
$\rho_{m}:X_{m}^{f}\rightarrow\bigsqcup_{j\in\mathcal{J}_{m}}\tilde{X}_{l_{j}},$ (61)
is straightforward, and we have $X^{f}_{m,j}=\rho_{m}^{-1}(\tilde{X}_{l_{j}})$ for all $j\in\mathcal{J}_{m}$. For $\tilde{D}\in\tilde{X}_{l_{j}}$, the inverse is given by $\tilde{D}\mapsto\pi^{*}(\tilde{D})+j.$ The proof that this is indeed an isomorphism uses that $p$ is a prime and is elementary. We will now show that $\delta:\mathcal{F}_{m}^{f}\rightarrow X_{m}^{f}$ is the projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle. The decomposition given in equation (61) will then entail that $\mathcal{M}_{i,j}^{\mathbb{T}}$, as defined in equation (59), is in fact a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle over $\tilde{X}_{l_{j}}$. The proof of the equivariant decomposition given in equation (60) will be given in the third step of the proof, and this will imply that $\mathcal{M}_{i,j}^{\mathbb{T}}$ is a subspace of $\mathcal{M}_{i}$. First we notice that $\gcd(p,2)=1$ implies that $\pi^{*}(\operatorname{Jac}(X))[2]=\operatorname{Jac}(X)[2]^{f}$, and it is elementary to verify that $\pi^{*}$ defines an isomorphism. Now notice that the image of $\operatorname{\mathcal{F}}_{m}^{f}$ under $\delta:\operatorname{\mathcal{F}}_{m}\rightarrow X_{m}$ belong to $X_{m}^{f}$. Further, if $\delta(\mathcal{E}_{1})=\delta(\mathcal{E}_{2})$ for $\operatorname{\mathcal{E}}_{1},\operatorname{\mathcal{E}}_{2}\in\operatorname{\mathcal{F}}_{m}^{f}$, then there exists a unique $L\in\operatorname{Jac}(X)[2]$ with $L.\mathcal{E}_{1}=\mathcal{E}_{2}$. Then
$f^{*}(L).\mathcal{E}_{1}$ $=f^{*}(L).f^{*}(\mathcal{E}_{1})=f^{*}(L.\mathcal{E}_{1})=f^{*}(\mathcal{E}_{2})=\mathcal{E}_{2}=L.\mathcal{E}_{1},$ (62)
and therefore the fact that $\operatorname{\mathcal{F}}_{m}$ is a $\operatorname{Jac}(X)[2]$-torsor implies that $f^{*}(L)=L$. Thus $L\in\operatorname{Jac}(X)[2]^{f}$. Therefore the Lemma will follow if we can establish surjectivity of the map $\delta:\operatorname{\mathcal{F}}_{m}^{f}\rightarrow X_{m}^{f}$. Given $\delta_{0}\in X_{m}^{f}$ we want to find $L_{0}\in\operatorname{Pic}(X)^{f}$, that solves the equation
$L_{0}^{2}=\Lambda K_{X}(-\delta_{0}).$ (63)
Recall that $K_{X}=\pi^{*}(K_{\tilde{X}})(-(p-1)X^{f})$. As $K_{\tilde{X}}$ is of even degree, it has a square-root, say $K_{\tilde{X}}^{1/2}$. As $p$ is odd, it follows that we have an $\langle f\rangle$-equivariant square-root of $K_{X}$ given by
$K_{X}^{1/2}=\pi^{*}(K_{\tilde{X}}^{1/2})(-2^{-1}(p-1)X^{f}).$ (64)
Thus we have reduced the problem to finding an $\langle f\rangle$-equivariant square-root of $\Lambda(-\delta_{0})$. Using the first part of the lemma, we see that there exists $j\in\mathcal{J}_{m}$ and $D\in\tilde{X}_{l_{j}}$, such that $\delta_{0}=\pi^{*}(D)+j$. As $p$ is an odd prime, we have $2\in(\mathbb{Z}/p\mathbb{Z})^{\times}$, and therefore we can for each $x\in X^{f}$ find $a_{x}\in\mathbb{Z}$ such that
$2a_{x}=-j_{x}\mod p.$
Define $A=\sum_{x\in X^{f}}a_{x}\cdot x.$ We get from Proposition 2.3 that $\mathcal{O}_{A}$ admits a canonical $\langle f\rangle$-equivariant structure, such that $\langle f\rangle$-acts as the identity on the restriction of $\mathcal{O}_{-(j+2A)}$ to $X^{f}.$ In particular, we can define the quotient line bundle
$L=\mathcal{O}_{-(j+2A)}/\langle f\rangle\in\operatorname{Pic}(\tilde{X}),$ (65)
and we have that $\mathcal{O}_{-(j+2A)}=\pi^{*}(L)$. Recall that $\Lambda=\pi^{*}(\tilde{\Lambda})$. We observe that
$\Lambda(-\delta_{0})=\Lambda(-\pi^{*}(D)-j-2A)\mathcal{O}_{2A}=\pi^{*}(\tilde{\Lambda}L(-D))\mathcal{O}_{2A}$ (66)
As $\mathcal{O}_{A}$ is an $\langle f\rangle$-equivariant square-root of $\mathcal{O}_{2A}$, we just need to find an $\langle f\rangle$-equivariant square-root of $\pi^{*}(\tilde{\Lambda}L(-D))$. As the degree of $\Lambda(-\delta_{0})$ is even, and $\deg(\pi^{*}(\tilde{\Lambda}L(-D))=p\deg\tilde{\Lambda}L(-D)$ we see that the degree of $\tilde{\Lambda}L(-D)$ must be even. Hence it admits a square-root, say $(\tilde{\Lambda}L(-D))^{1/2}$, and therefore $\pi^{*}((\tilde{\Lambda}L(-D))^{1/2})$ is an $\langle f\rangle$-equivariant square-root of $\pi^{*}(\tilde{\Lambda}L(-D))$. This finishes the proof of the first part.
$2$. Let $x\in X^{f}$. As $f_{\operatorname{\mathbbm{E}}}$ commutes with the $\mathbb{T}$-action, and as $\mathbb{T}$ act with weight $t$ on $\operatorname{\mathbbm{E}}_{m,1}$ and with weight equal to unity on $\operatorname{\mathbbm{E}}_{m,0}$, it easily follows that we must have a $\mathbb{T}\times\langle f\rangle$-equivariant decomposition of the form (60), for some $\mu^{\prime}_{x}\in\mathbb{Z}/p\mathbb{Z}$ in place of $\epsilon_{j,x}i_{x}$. Thus we must prove that
$2\mu^{\prime}_{x}=-n_{x}(1+j_{x})\mod p.$ (67)
Let $\operatorname{\mathcal{E}}=(L_{1}\oplus L_{2},\Phi_{b})\in\mathcal{M}_{i,j}^{\mathbb{T}}$ be given. Consider a frame $(s_{1},s_{2})$ near $x$, where $s_{j}$ is a frame of $L_{j}$ for $j=1,2$. Let $z$ be a coordinate centered at $x$. Write $b=\phi(s_{2}\otimes s_{1}^{*})dz$, where $\phi\in\mathcal{O}_{X,x}$ is a regular function vanishing to order $d$ at $x$. Then $d=j_{x}\mod p$ by definition. Thus we may write $\phi=z^{d}\cdot h_{1}(z)$ with $h_{1}(x)\not=0.$ It follows that $f^{*}(\phi)=\zeta^{n_{x}j_{x}}z^{d}f^{*}(h_{1})$. As $df^{*}(f^{*}(dz))=\frac{\partial f}{\partial z}dz=\zeta^{n_{x}}dz$, we obtain
$(\operatorname{Id}\otimes df^{*})(f^{*}(\Phi_{b}))=\begin{pmatrix}0&0\\ \zeta^{n_{x}(j_{x}+1)}z^{d}f^{*}(h_{1})&0\end{pmatrix}dz$ (68)
On the other hand, we may write $f_{\operatorname{\mathcal{E}}}$ with respect to the frames $(s_{1},s_{2})$ and $(f^{*}(s_{1}),f^{*}(s_{2}))$ as a diagonal matrix $f_{\operatorname{\mathcal{E}}}=\operatorname{diag}(f_{1},f_{2})$, where $f_{1},f_{2}\in\mathcal{O}_{X,x}$ are regular functions defined near $x$ with $f_{1}(x)=\sqrt{\alpha}_{x}\zeta^{\mu^{\prime}_{x}}$ and $f_{2}(x)=\sqrt{\alpha}_{x}\zeta^{-\mu^{\prime}_{x}}$. Recall that $f^{c}_{\operatorname{\mathcal{E}}}$ denotes conjugation by $f_{\operatorname{\mathcal{E}}}$. Then we have that
$\begin{split}&(f^{c}_{\operatorname{\mathcal{E}}}\otimes\operatorname{Id}_{K})(\Phi_{b})=\begin{pmatrix}0&0\\ f_{2}f_{1}^{-1}\phi&0\end{pmatrix}dz.\end{split}$ (69)
As $\operatorname{\mathcal{E}}$ is a fixed point, equation (19) holds for $\Phi_{b}$, and thus the right hand side of equation (68) and the right hand side of equation (69) are equal, i.e. we have that
$\zeta^{n_{x}(j_{x}+1)}z^{d}f^{*}(h_{1})dz=f_{2}f_{1}^{-1}z^{d}h_{1}dz.$ (70)
Dividing out by $z^{d}$ on both sides of equation (70), evaluating at $z=0$ and using $h_{1}(0)\not=0$, we see that equation (67) holds. ∎
Fix $i\in\mathcal{I}.$ Towards the end of identifying the components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$, we note that $\tilde{\mathcal{N}}_{i}$ is always one of the components, so it suffices to identify the components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}\setminus\tilde{\mathcal{N}}_{i}$. By the above lemma this is equivalent to identifying the image $\varXi_{i}(\mathcal{M}_{i,j}^{\mathbb{T}})$ for all $m\in M$ and $j\in\mathcal{J}_{i,m}.$
Definition 2.8.
Set $\mathcal{J}^{\prime}_{i}=\sqcup_{m\in M}\mathcal{J}_{i,m}$. For $j\in\mathcal{J}^{\prime}_{i}$ denote by ~$\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ the set of isomorphism classes of quasi parabolic Higgs bundles $(E,F,\Phi_{b})$ on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ of rank two of the following type.
  • There are $L_{+},L_{-}\in\operatorname{Pic}(\tilde{X})$ with $E=L_{+}\oplus L_{-}$ and $L_{+}L_{-}\simeq\tilde{\Lambda}_{i}$, and for each $\tilde{x}\in\operatorname{\tilde{\mathcal{D}}}_{i}$, we have that $F(\tilde{x})=L_{\epsilon_{j,x}}(\tilde{x})$.
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(\tilde{X},L_{+}L_{-}^{-1}K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}))$. The zero divisor of $b$ has degree $l_{j}$.
Corollary 2.1.
Fix $i\in\mathcal{I}$. Set $\mathcal{J}_{i}=\{i\}\sqcup\mathcal{J}_{i}^{\prime}$ and set ~$\tilde{\mathcal{M}}^{\mathbb{T}}_{i,i}=\tilde{\mathcal{N}}_{i}$. Fix $j\in\mathcal{J}^{\prime}_{i}$. Then $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. Thus the set of components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$ is in bijection with $\mathcal{J}_{i}$. In particular, every ~$\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with resepct to the weight $w_{i}$.
Proof.
The proof follows from an elementary analysis of how the Hecke modification $\Psi_{i}$ act on $\operatorname{\mathcal{E}}=(L_{1}\oplus L_{2},\Phi_{b})\in\mathcal{M}_{i,j}^{\mathbb{T}}$. Consider the setup of the second step of the proof of Proposition 2.4. Let $x\in\mathcal{D}_{i}$. Assume for definiteness that $\epsilon_{j,x}=1$. We can choose the frame $(s_{1},s_{2})$ of $\operatorname{\mathcal{E}}$ such that $s_{1}$ is a non-vanishing section of $L_{1}$ and $s_{2}$ is a non-vanishing section of $L_{2}$. Then the Hecke transform is given as follows
$\operatorname{\mathcal{E}}^{\prime}=(L_{1}\oplus L_{2}(-x),\Phi_{b}^{\prime}).$ (71)
Write $E=L_{1}\oplus L_{2}$ and $E^{\prime}=L_{1}\oplus L_{2}(-x)$. Notice that in the frame $(s_{1},s_{2})$, the matrix of $f_{E}$ is diagonal in a neighbourhood of $x$. Therefore equation (30) implies that the lift $f_{E^{\prime}}$ of the Hecke transform $\operatorname{\mathcal{E}}^{\prime}$ is diagonal with respect to any frame of the form $(s_{1}^{\prime},s_{2}^{\prime})$ where $s_{1}^{\prime}$ is a non-vanishing section of $L_{1}$ and $s_{2}^{\prime}$ is a non-vanishing section of $L_{2}(-x)$. Similarly, equation (27) implies that $\Phi^{\prime}_{b}$ will also be an off-diagonal Higgs field associated with a section.
That every ~$\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with respect to $w_{i}$ follows from Theorem 1.1, the fact that $\mathcal{M}^{\mathbb{T}}_{i,j}$ is a sub-variety of $\mathcal{M}_{i}$ and the identity $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. ∎
Remark 2.
From Corollary 2.1 it follows that every connected component is either a smooth moduli space of stable parabolic vector bundles on $\tilde{X}$ of rank two, or a principal-$\operatorname{Jac}(\tilde{X})[2]$ bundle over a symmetric power of $\tilde{X}$. Consider first the case of moduli spaces of stable bundles. The intersection paring of these moduli spaces have been thoroughly studied following Witten’s work [66]. In the so-called coprime case, formulae for all pairings were proven in rank two in [59] and in all ranks in [42]. Subsequently, a version of these formulae for a class of parabolic moduli spaces including $\tilde{\mathcal{N}}_{i}$ were proven in [meinrenkenWittenFormulaeIntersection2005]. Consider the case of the intersection pairings on the cohomology ring of symmetric powers of a Riemann surface. These rings are generated by MacDonald classes [49] and the intersection pairings were computed by Zagier, and the resulting formulae can be found in [60].
3 Localization
We collect some generalities on localization before proving Theorem 1.2. Let $H$ be a finite cyclic group with generator $h.$ Let $(X,\omega)$ be a Kähler manifold with a holomorphic $\mathbb{T}\times H$-action, for which the $\mathbb{T}$-action is meromorphic in the sense of [68], i.e. assume Assumption $2.15$ in [68, §2.2] is satisfied. Let $L\rightarrow X$ be a holomorphic $\mathbb{T}\times H$-equivariant line bundle of curvature $-i\omega$. Then $L\rightarrow X$ is a quantum line bundle for $(X,\omega)$ in the sense of geometric quantization [47, 58]. For all $q\in\mathbb{Z}_{+}$ and $m\in\mathbb{Z}$ denote by $H^{q}_{m}(X,L)$ the $t^{m}$-weight sub-space of $H^{q}(X,L)$. Define
$\chi_{\mathbb{T}}(X,L,h)(t)=\sum_{m\in\mathbb{Z}}\sum_{q\geq 0}(-1)^{q}\operatorname{tr}(h:H_{m}^{q}(X,L)\rightarrow H_{m}^{q}(X,L))\cdot t^{m}.$ (72)
The meromorphicity condition on the $\mathbb{T}$-action on $X$ implies that the connected components of $X^{\mathbb{T}}$ are compact smooth complex manifolds. Write $A:=\pi_{0}(X^{\mathbb{T}\times H})$ and denote by $\{X_{a}\}_{a\in A}$ the components of $X^{\mathbb{T}\times H}$. Fix $a\in A$. Let $T_{a}X$ denote the restriction of $TX$ to $X_{a}$. For every $\mathbb{T}\times H$-equivariant vector bundle $W\rightarrow X_{a}$ denote by $W^{\pm}$ the $\pm$-$\mathbb{T}$-weight subbundle of $W$, and define $\nu_{a}=\operatorname{rank}(T_{a}X^{-}).$ Recall the notation $s$ for the equivariant sum of all symmetric powers. Define the equivariant class
$\omega(W)=s((W^{+})^{*})\cdot s(W^{-})\cdot\det(W^{-}).$ (73)
Denote by $N^{*}(X_{a})$ the conormal of $X_{a}$ as a submanifold of $X^{\mathbb{T}}$. Denote by $\lambda=\sum_{l\geq 0}(-1)^{l}\lambda^{l}$ the equivariant operation of taking the alternating sum of exterior powers.
Proposition 3.1.
We have that
$\begin{split}&\chi_{\mathbb{T}}(X,L,h)=\sum_{a\in A}(-1)^{\nu_{a}}\int_{X_{a}}\operatorname{ch}\left(\frac{\omega(T_{a}X)\cdot L_{\mid X_{a}}}{\lambda\left(N^{*}(X_{a})\right)}\right)\operatorname{Td}X_{a}.\end{split}$ (74)
Proposition 3.1 follows essentially from a combination of a $\mathbb{T}$-localization result of Wu [68] with the Atiyah-Bott fixed point formula [14], also known as the Lefshetz fixed-point formula. In [68], Wu considers the case $H=\langle 1\rangle$. In this case $\{X_{a}\}_{a\in A}$ is equal to the set of components of $X^{\mathbb{T}}$ and in [68, §4.2] the following is proven.
Theorem 3.1 ([68]).
In $\mathbb{T}$-equivariant $K$-theory we have
$\begin{split}&\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}(-1)^{q}H^{q}_{m}(X,L)\cdot t^{m}=\sum_{\begin{subarray}{c}m\in\mathbb{Z},\\ q\geq 0\end{subarray}}\sum_{a\in A}(-1)^{\nu_{a}+q}H_{m}^{q}(X_{a},\omega(T_{a}X)\cdot L_{\mid X_{a}})\cdot t^{m}.\end{split}$ (75)
We now recall the Atiyah-Bott Lefshetz formula [14]. Let $Z$ be a compact complex manifold with a holomorphic $H$-action. Let $W\rightarrow Z$ be an $H$-equivariant holomorphic vector bundle. Consider the equivariant index
$\chi(Z,W,h)=\sum_{q\geq 0}(-1)^{q}\operatorname{tr}(h:H^{q}(Z,W)\rightarrow H^{q}(Z,W)).$
Let $Z^{H}$ denote the fixed locus. Write $B:=\pi_{0}(Z^{H})$ and denote by $\{Z_{b}\}_{b\in B}$ the components of $Z^{H}$. For each component $Z_{b}$, denote by $N^{*}(Z_{b})$ the conormal of $Z_{b}\subset Z$. The components are of course compact and smooth.
Theorem 3.2 ([14]).
We have that
$\begin{split}&\chi(Z,W,h)=\sum_{b\in B}\int\limits_{Z_{b}}\operatorname{ch}\left(\frac{W_{\mid Z_{b}}}{\lambda(N^{*}(Z_{b}))}\right)\operatorname{Td}Z_{b}.\end{split}$ (76)
Proof of Proposition 3.1.
The proof is a simple computation which combines (75) with (76). The key point is that the identity (75) can be generalized to an identity in the $K$-theory ring of $\mathbb{T}\times H$-representations. ∎
We now recall Nielsens’ localization result [52]. Assume now that $H=\mu_{p}$ and that $Z$ is a smooth projective $\mu_{p}$-variety. Let $K_{\mu_{p}}^{0}(Z)$ denote the $K$-theory ring of $\mu_{p}$-equivariant holomorphic vector bundles. Let $Z^{\mu_{p}}$ denote the fixed variety. Recall that $\mathbb{Z}[\mu_{p}]$ denotes the representation ring of $\mu_{p}$. Recall the multiplicatively closed subset $S\subset\mathbb{Z}[\mu_{p}]$ defined in Section 1.2. We can regard $S$ as a multiplicatively closed subset of $K^{0}_{\mu_{p}}(Z)$ and of $K^{0}_{\mu_{p}}(Z^{\mu_{p}})$.
Theorem 3.3 ([52]).
Pull-back along the inclusion $i:Z^{\mu_{p}}\rightarrow Z$ induces an isomorphism $S^{-1}\cdot K_{\mu_{p}}^{0}(Z)\overset{\sim}{\rightarrow}S^{-1}\cdot K_{\mu_{p}}^{0}(Z^{\mu_{p}})$. The inverse is given by
$W\mapsto i_{!}(W\cdot\lambda^{-1}(N^{*}(Z^{\mu_{p}}))).$ (77)
We now analyze $K_{\mu_{p}}^{0}(Z^{\mu_{p}})$ in more detail. We have canonical isomorphisms
$K_{\mu_{p}}^{0}(Z^{\mu_{p}})\cong K^{0}(Z^{\mu_{p}})\bigotimes_{\mathbb{Z}}\mathbb{Z}[\mu_{p}]\cong\bigoplus_{j=0}^{p-1}\operatorname{K}^{0}(Z^{\mu_{p}})\cdot\hat{\zeta}^{j}.$ (78)
For $j\in\{0,...,p-1\}$ write $\pi_{j}$ for the projection onto the $j$’th summand in (78). We introduce the special notation $W_{0}=\pi_{0}(W).$ We denote the universal localization homomorphism by $L:K^{0}_{\mu_{p}}(Z^{\mu_{p}})\rightarrow S^{-1}\cdot K^{0}_{\mu_{p}}(Z^{\mu_{p}}).$ Recall that the kernel of $L$ is generated by all elements $r$ such that $s\cdot r=0$ for some $s\in S$. Let $Y$ be a formal variable. Recall that $\mathbb{Z}[\mu_{p}]\cong\mathbb{Z}[Y]/(Y^{P}-1)$ and that $Y^{P}-1=(Y-1)\Phi_{p}(Y)$ is the factorization of $Y^{p}-1$ into irreducible factors. As $(1-\hat{\zeta})\in S$, it is therefore clear that we have $\ker(L)=K^{0}(Z^{\mu_{p}})\otimes_{\mathbb{Z}}(\Phi_{p}(\hat{\zeta})).$ Moreover, the following sequence of abelian groups is split exact
[…]
and a complement to $\ker(L)$ is given by the module generated by all classes of the form $W-W_{0}\cdot\Phi_{p}(\hat{\zeta})$, where $W$ ranges over all equivariant bundles. The following lemma is an elementary consequence of these facts.
Lemma 3.1.
If $V,W\in K_{\mu_{p}}^{0}(Z^{\mu_{p}})$ are classes such that $V=W\mod\ker(L)$, then the following identity holds in $K_{\mu_{p}}^{0}(Z^{\mu_{p}})$
$V-V_{0}\cdot\Phi_{p}(\hat{\zeta})=W-W_{0}\cdot\Phi_{p}(\hat{\zeta}).$ (80)
Recall now the setup and the notation introduced in Section 1.2. For each $l\in\{0,...,p-1\}$, let $\pi_{l}$ be the projection onto the $l$’th summand of $R_{i,j}$ in (5).
Definition 3.1.
For all $i\in\mathcal{I}$, all $j\in\tilde{\mathcal{J}_{i}}$ and all $l\in\{1,...,p-1\}$ define the integer $\tilde{\nu}_{i,j}$ and the rationals $\tilde{a}_{i,j,l}$ and $\tilde{b}_{i,j,l}$ by the formulae
$\tilde{\nu}_{i,j}=\operatorname{rank}(T_{i,j}\tilde{\mathcal{M}}_{i}^{-})+\operatorname{ch}_{0}(\tilde{V}_{i,j}^{-}-\pi_{0}(\tilde{V}_{i,j}^{-})),$ (81)
$\tilde{a}_{i,j,l}=\operatorname{rank}(T\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j})+\frac{\operatorname{ch}_{0}\left(\pi_{l}(\tilde{V}^{\mathbb{T}}_{i,j})-\pi_{0}(\tilde{V}^{\mathbb{T}}_{i,j})\right)}{p},$ (82)
$\tilde{b}_{i,j,l}=\operatorname{rank}(T_{i,j}\tilde{\mathcal{M}}_{i}^{-})+\frac{\operatorname{ch}_{0}\left(\pi_{l}(\tilde{V}^{-}_{i,j})-\pi_{0}(\tilde{V}^{-}_{i,j})\right)}{p}.$ (83)
Set $\tilde{\rho}_{i,j}=\sum_{l=1}^{p-1}l\cdot\tilde{b}_{i,j,l}$ and define $\tilde{\zeta}_{i,j}=\hat{\zeta}^{\tilde{\rho}_{i,j}}\prod_{l=1}^{p-1}(1-\hat{\zeta}^{l})^{-\tilde{a}_{i,j,l}}.$
It will be seen in the proof of Theorem 1.2 that the $\tilde{a}_{i,j,l}$ and the $\tilde{b}_{i,j,l}$ are in fact positive integers. Before given the proof of Theorem 1.2 we state the following Lemma.
Lemma 3.2.
Let $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$. Set $d=\dim_{\mathbbm{C}}(\mathcal{M}_{i,j}^{\mathbb{T}})$ and define the ring $R=\oplus_{l=0}^{d}H^{2l}(\mathcal{M}_{i,j}^{\mathbb{T}},\mathbbm{Q})\otimes_{\mathbb{Z}}(S^{-1}\cdot\mathbb{Z}[\mu_{p}])$. Assume $a,b\in R[[t^{-1}]]$ satisfy $a^{p}=b^{p}$, and that the degree zero components (with respect to the cohomological-grading) of the leading terms (with respect to the $t^{-1}$-grading) of $a$ and $b$ are equal and non-zero. Then $a=b$.
Proof.
Write $a=\sum_{i=0}^{\infty}a_{i}t^{-1}$ and $b=\sum_{i=0}^{\infty}b_{i}t^{-i}$ with $a_{i},b_{i}\in R$ for all non-negative integers $i$. Without loss of generality, we may assume $a_{0}\not=0$ and $b_{0}\not=0$. We have
$\begin{split}&a^{p}=\sum_{i=0}^{\infty}\sum_{j_{1}+\cdots+j_{p}=i}a_{j_{1}}\cdots a_{j_{p}}t^{-i}=b^{p}=\sum_{i=0}^{\infty}\sum_{j_{1}+\cdots+j_{p}=i}b_{j_{1}}\cdots b_{j_{p}}t^{-i}.\end{split}$ (84)
In particular, we have that $a_{0}^{p}=b_{0}^{p}$. For each positive integer $l$, let $R^{l}$ denote the degree $2l$ part of $R$ with respect to the cohomological grading. Write $a_{0}=\sum_{l=0}^{d}a_{0,l}$ and $b_{0}=\sum_{l=0}^{d}b_{0,l}$ with $a_{0,l},b_{0,l}\in R^{l}$ for $l=0,...,d$. Then we have $a_{0,0}=b_{0,0}$ by assumption, and we see that
$\begin{split}&a_{0}^{p}=\sum_{l=0}^{d}\sum_{l_{1}+\cdots+l_{p}=l}a_{0,l_{1}}\cdots a_{0,l_{p}}=b_{0}^{p}=\sum_{i=0}^{d}\sum_{l_{1}+\cdots+l_{p}=l}b_{0,l_{1}}\cdots b_{0,l_{p}}.\end{split}$ (85)
For each $l=1,..,d$ we get by taking the degree $l$ part of (85) that
$\begin{split}&\sum_{l_{1}+\cdots+l_{p}=l}a_{0,l_{1}}\cdots a_{0,l_{p}}=pa_{0,0}^{p-1}a_{0,l}+\sum_{\begin{subarray}{c}l_{1}+\cdots+l_{p}=l,\\ l_{t}<l,t=1,...,p\end{subarray}}a_{0,l_{1}}\cdots a_{0,l_{p}}\\ &=\sum_{l_{1}+\cdots+l_{p}=l}b_{0,l_{1}}\cdots b_{0,l_{p}}=pb_{0,0}^{p-1}b_{0,l}+\sum_{\begin{subarray}{c}l_{1}+\cdots+l_{p}=l,\\ l_{t}<l,t=1,...,p\end{subarray}}b_{0,l_{1}}\cdots b_{0,l_{p}}.\end{split}$ (86)
Now notice that any element of $R$ whose projection onto $R^{0}$ is non-zero is not a zero divisor in $R$. In particular $a_{0,0}^{p}=b_{0,0}^{p}$ is not a zero divisor, and therefore we get by induction from (86) that $a_{0,l}=b_{0,l}$ for $l=0,...,d$. Thus $a_{0}=b_{0}$. From (84) we deduce that for all non-negative integers $i$ we have that
$\begin{split}&\sum_{j_{1}+\cdots+j_{p}=i}a_{j_{1}}\cdots a_{j_{p}}=pa_{0}^{p-1}a_{i}+\sum_{\begin{subarray}{c}j_{1}+\cdots+j_{p}=i,\\ j_{t}<i,t=1,...,p\end{subarray}}a_{j_{1}}\cdots a_{j_{p}}\\ &=\sum_{j_{1}+\cdots+j_{p}=i}b_{j_{1}}\cdots b_{j_{p}}=pb_{0}^{p-1}b_{i}+\sum_{\begin{subarray}{c}j_{1}+\cdots+j_{p}=i,\\ j_{t}<i,t=1,...,p\end{subarray}}b_{j_{1}}\cdots b_{j_{p}}.\end{split}$ (87)
By induction, and the fact that $a_{0}^{p}$ and $b_{0}^{p}$ are not zero divisors, we obtain from (87) that $a_{i}=b_{i}$ for all non-negative integers $i$. Thus $a=b$. ∎
Proof of Theorem 1.2.
Fix $i\in\mathcal{I}$. Write $\alpha_{i}$ (resp. $\tilde{\alpha}_{i}$) for the equivariant first Chern class of the restriction of $\mathcal{L}$ (resp. ~$\tilde{\mathcal{L}}_{i}$) to $\mathcal{M}_{i}$ (resp. ~$\tilde{\mathcal{M}}_{i}$). Then we have that
$p\cdot\tilde{\alpha}_{i}=(\varXi_{i}^{-1})^{*}(\alpha_{i}).$ (88)
For the restriction $\varXi_{i}:\mathcal{N}\cap\mathcal{M}_{i}\rightarrow\tilde{\mathcal{N}}_{i}$, the identity (88) was in fact proven in the article [3], and the proof therein can be generalized to the isomorphism of $\mathbb{T}$-varieties $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$. We now turn to the proof of equation (9). The bundle $\mathcal{L}\rightarrow\mathcal{M}$ satisfies the conditions of Proposition 3.1 with $H=\langle f\rangle$ and the $\mathbb{T}\times\langle f\rangle$-equivariant structure constructed in Section 2. Thus we can apply Proposition 3.1. Let $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$. To ease notation, we will write $\mathcal{M}_{i,j}^{\mathbb{T}}=\mathcal{M}_{i,j}$ and we will write $T_{i,j}\mathcal{M}$ (resp. $T_{i,j}\mathcal{M}_{i}$) for the restriction of $T\mathcal{M}$ (resp. $T\mathcal{M}_{i}$) to $\mathcal{M}_{i,j}$. Let $N^{*}_{i,j}$ be the conormal of $\mathcal{M}_{i,j}\subset\mathcal{M}^{\mathbb{T}}$. By Proposition 3.1, it will be sufficient to prove that for all $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$ we have that
$(-1)^{\nu_{i,j}}\omega(T_{i,j}\mathcal{M})\cdot\lambda\left(N^{*}_{i,j}\right)^{-1}=(-1)^{\tilde{\nu}_{i,j}}\varXi_{i}^{*}(\tilde{\chi}_{i,j}).$ (89)
Write $\mathbb{U}=\operatorname{End}_{0}(\operatorname{\mathbbm{E}})$, and define the equivariant class
$V_{i,j}=\sum_{x\in X^{f}}\mathbb{U}_{\mid\mathcal{M}_{i,j}\times\{x\}}\cdot(1-t\cdot\hat{\zeta}^{-n_{x}})\sum_{l=1}^{p-1}l\cdot\hat{\zeta}^{-ln_{x}}.$ (90)
Then $\varXi_{i}^{*}(\tilde{V}_{i,j})=V_{i,j}$. Let $\pi:\mathcal{M}^{\mathbb{T}}\times X\rightarrow\mathcal{M}^{\mathbb{T}}$ be the projection onto the first factor. On $\mathcal{M}_{i,j}\subset\mathcal{M}^{\mathbb{T}}$, we have the following equation in $\mathbb{T}$-equivariant $K$-theory
$T_{i,j}\mathcal{M}=\pi_{!}(\mathbb{U}\cdot\pi^{*}(K_{X})\cdot t-\mathbb{U}).$ (91)
This can be seen for instance by considering a standard long exact sequence for hypercohomology groups of two-term chain complexes (see [36, §2.2.3]). Consider the natural projection onto the first factor $\pi^{\prime}:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}$. Note that we have $(\mathcal{M}_{i,j}\times X)^{f}=\mathcal{M}_{i,j}\times X^{f}$. Consider the inclusion $\iota:\mathcal{M}_{i,j}\times X^{f}\rightarrow\mathcal{M}_{i,j}\times X$. Clearly $\pi\circ\iota=\pi^{\prime}.$ By applying Nielsen’s localization theorem to the right hand side of quation (91), we obtain the following identity in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$
$T_{i,j}\mathcal{M}$ $=\pi_{!}\left(\iota_{!}\left(\frac{\iota^{*}(\mathbb{U}(\pi^{*}(K_{X})\cdot t-1))}{\lambda(N^{*}(\mathcal{M}_{i,j}\times X^{f}))}\right)\right)=\pi^{\prime}_{!}\left(\frac{\iota^{*}(\mathbb{U}(\pi^{*}(K_{X})\cdot t-1))}{\lambda(N^{*}(\mathcal{M}_{i,j}\times X^{f}))}\right)$ (92)
$=\sum_{x\in X^{f}}\frac{\mathbb{U}_{\mid\mathcal{M}_{i,j}\times\{x\}}\cdot(1-t\cdot\hat{\zeta}^{-n_{x}})}{\hat{\zeta}^{-n_{x}}-1}.$ (93)
However, the $\mathbb{T}\times\mu_{p}$-equivariant $K$-theory operation $s$ of taking the sum of all symmetric powers is not well-defined in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\mu_{p}}(\mathcal{M}_{i,j})$, and we need to extract an identity in the ring $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})$. For all $u=1,...,p-1$, we have the following identity in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$
$p\cdot(\hat{\zeta}^{u}-1)^{-1}=\sum_{l=1}^{p-1}l\hat{\zeta}^{ul},$ (94)
which easily follows from the fact that $\Phi_{p}(\hat{\zeta}^{u})=0$ in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$ for $u=1,...,p-1$. Multiplying by $p$ and using equation (94) we obtain the following identity in $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})/\ker(L)$
$p\cdot T_{i,j}\mathcal{M}=V_{i,j}\mod\ker(L).$ (95)
Notice that $\pi_{0}(T_{i,j}\mathcal{M})=T_{i,j}\mathcal{M}_{i}$. Thus equation (95) together with Lemma 3.1 implies that the following identity holds in the ring $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}^{\mathbb{T}}_{i,j})$
$p\cdot T_{i,j}\mathcal{M}=p\cdot T_{i,j}\mathcal{M}_{i}\cdot\Phi_{p}(\hat{\zeta})+V_{i,j}-\pi_{0}(V_{i,j})\cdot\Phi_{p}(\hat{\zeta}).$ (96)
We can now apply $\omega$ to equation (96). Using that $\omega$ is a homomorphism that transforms sums into products, we obtain
$\omega(T_{i,j}\mathcal{M})^{p}=\omega(T_{i,j}\mathcal{M}_{i}\cdot\Phi_{p}(\hat{\zeta}))^{p}\frac{\omega(V_{i,j})}{\omega(\pi_{0}(V_{i,j})\cdot\Phi_{p}(\hat{\zeta}))}.$ (97)
Let $u$ be a formal variable. Denote by $s_{u}=\sum_{m\geq 0}u^{m}s^{m}$ the operation of taking the graded sum of symmetric powers, and denote by $\lambda_{u}=\sum_{m\geq 0}u^{m}\lambda^{m}$ the operation of taking the graded sum of exterior powers. Recall that for any equivariant vector bundle $V$ we have that $s_{u}(V)\lambda_{-u}(V)=1$. Therefore we introduce the notation $\lambda(W)^{-1}=s(W)$ for any equivariant bundle $W$ with trivial $\mathbb{T}$-action for which this class exists in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\langle f\rangle}(\mathcal{M}_{i,j})$. By Nielsen’s localization theorem (Theorem 3.3), this holds for $N^{*}_{i,j}$. We have
$s\left(N^{*}_{i,j}\right)=\lambda(T^{*}\mathcal{M}_{i,j})\cdot s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}}).$ (98)
Now we will rewrite the factor $s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}}).$ Using that $\pi_{0}(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}})=T^{*}\mathcal{M}_{i,j}$ and the homomorphism property of $s$, we obtain from (96) the following equation
$\begin{split}s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}})^{p}&=s(T^{*}\mathcal{M}_{i,j}\cdot\Phi_{p}(\hat{\zeta}))^{p}\prod_{x\in X^{f}}\frac{s((V^{\mathbb{T}}_{x,i,j})^{*})}{s(\pi_{0}(V^{\mathbb{T}}_{x,i,j})^{*}\cdot\Phi_{p}(\hat{\zeta}))}.\end{split}$ (99)
We now observe that for all equivariant vector bundles $W\rightarrow\mathcal{M}_{i,j}$, for which $s(W)$ exists, we have that
$\Omega(W)=\omega(W)\cdot s((W^{\mathbb{T}}_{i,j})^{*}).$ (100)
By multiplying by $\lambda(N^{*}_{i,j})^{-p}$ on both sides of equation (97), and using equation (98), equation (99), equation (100) and the definition of $\tilde{\chi}_{i,j}$ given in equation (8), we obtain the following identity
$\omega(T_{i,j}\mathcal{M})^{p}\cdot\lambda\left(N^{*}_{i,j}\right)^{-p}=\varXi_{i}^{*}(\tilde{\chi}_{i,j})^{p}.$ (101)
Since $p$ is odd, it is trivial to see that equation (96) implies that $(-1)^{\nu_{i,j}}=(-1)^{\tilde{\nu}_{i,j}}.$ Thus it only remains to extract the correct $p$’th root. For each $l=1,...,p-1$ define $a_{i,j,l}=\operatorname{rank}(\pi_{l}(N^{*}_{i,j}))$, and define $b_{i,j,l}=\operatorname{rank}(\pi_{l}(T_{i,j}\mathcal{M}^{-}))$. Further, define $\rho_{i,j}=\sum_{l=1}^{p-1}b_{i,j,l}\cdot l.$ Consider the projection, which takes $\operatorname{ch}_{0}$ of the leading term of a $1/t$-series
$P_{0}:H^{*}(\mathcal{M}_{i,j},\mathbbm{Q})\otimes\mathbbm{Q}[\mu_{p}][[t^{-1}]]\rightarrow H^{0}(\mathcal{M}_{i,j},\mathbbm{Q})\otimes(S^{-1}\cdot\mathbb{Z}[\mu_{p}]).$ (102)
Upon inspecting the definition of $\omega$ one finds that
$P_{0}\left(\operatorname{ch}(\omega(T_{i,j}\mathcal{M})\cdot\lambda\left(N^{*}_{i,j})\right)^{-1}\right)=\hat{\zeta}^{\rho_{i,j}}\prod_{l=1}^{p-1}(1-\hat{\zeta}^{l})^{-a_{i,j,l}}.$ (103)
From equation (96) we deduce that
$p\left(T_{i,j}\mathcal{M}^{\mathbb{T}}-T\mathcal{M}_{i,j}\right)=p\sum_{l=1}^{p-1}T\mathcal{M}_{i,j}\cdot\hat{\zeta}^{l}+V^{\mathbb{T}}_{i,j}-\pi_{0}(V_{i,j}^{\mathbb{T}})\cdot\Phi_{p}(\hat{\zeta}),$ (104)
$pT_{i,j}\mathcal{M}^{-}=p(T_{i,j}\mathcal{M}_{i})^{-}\cdot\Phi_{p}(\hat{\zeta})+V_{i,j}^{-}-\pi_{0}(V_{i,j})^{-}\cdot\Phi_{p}(\hat{\zeta}),$ (105)
and therefore, we obtain
$a_{i,j,l}=\dim(\mathcal{M}^{\mathbb{T}}_{i,j})+\frac{\operatorname{ch}_{0}\left(\pi_{l}(V_{i,j}^{\mathbb{T}})-\pi_{0}({V_{i,j}^{\mathbb{T}})}\right)}{p},$ (106)
$b_{i,j,l}=\operatorname{rank}(T_{i,j}\mathcal{M}_{i}^{-})+\frac{\operatorname{ch}_{0}\left(\pi_{l}(V_{i,j}^{-})-\pi_{0}(V_{i,j}^{-})\right)}{p}.$ (107)
Thus it is clear that $a_{i,j,l}=\tilde{a}_{i,j,l}$ and $b_{i,j,l}=\tilde{b}_{i,j,l}$. Therefore also $\rho_{i,j}=\tilde{\rho}_{i,j}$. It follows that the right hand side of equation (103) is equal to $\tilde{\zeta}_{i,j}$. Thus we can conclude by applying Lemma 3.2 applied to the elements $a=\operatorname{ch}(\tilde{\chi}_{i,j})$ and $b=\operatorname{ch}(\omega(T_{i,j}\mathcal{M})\cdot\lambda(N^{*}_{i,j})^{-1}).$
3.1 The Galois Case
Recall the notation $\pi_{\mathcal{M}}:\mathcal{M}\times X\rightarrow\mathcal{M}$ for the natural projection onto the first factor. Assume that $X^{f}=\emptyset$. Then Hurwitz formula specialises to $\chi(X)=p\chi(\tilde{X})$, where $\chi(X)$ (resp. $\chi(\tilde{X}))$) denotes the Euler characteristic of $X$ (resp. $\tilde{X}$). Fix $\tilde{x}_{0}\in\tilde{X}$ and define
$\mathcal{L}=\det((\pi_{\mathcal{M}})_{!}(\operatorname{\mathbbm{E}})[1])\bigotimes_{x\in\pi^{-1}(\tilde{x}_{0})}\det(\operatorname{\mathbbm{E}}_{\mid\mathcal{M}\times\{x\}})^{1-g}.$ (108)
Proof of Corollary 1.2.
In the case $X^{f}$ is empty, pullback with respect to the projection $\pi:X\rightarrow\tilde{X}$ induces an isomorphism of $\mathbb{T}$-varities $\pi^{*}:\tilde{\mathcal{M}}\rightarrow\mathcal{M}^{f}$. We write $\varXi=(\pi^{*})^{-1}$ for the inverse isomorphism of $\mathbb{T}$-varieties. We have that $(\varXi^{-1})^{*}(\mathcal{L})\cong\tilde{\mathcal{L}}^{\otimes p}$. We simplify the notation and write $\{\tilde{F}_{a}\}_{a\in A}$ for the components of ~$\tilde{\mathcal{M}}^{\mathbb{T}}$. Let $a\in A$. Denote by $T_{a}\tilde{\mathcal{M}}$ the restriction of $T\tilde{\mathcal{M}}$ to $\tilde{F}_{a}$ and write ${\nu}_{a}=\operatorname{rank}(T_{a}\tilde{\mathcal{M}}^{-})$. Observe that as $\operatorname{\tilde{\mathcal{D}}}$ is empty, we have that $\tilde{\nu}_{a}=\nu_{a}$. Define the $t^{-1}$-series
$\chi_{a}(t)=(-1)^{\tilde{\nu}_{a}}\int_{\tilde{F}_{a}}\operatorname{ch}(\omega(T_{a}\tilde{\mathcal{M}})\cdot\tilde{\mathcal{L}}^{k})\cdot\operatorname{Td}\tilde{F}_{a}.$ (109)
Then $\chi_{t}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})(t)=\sum_{a\in A}\chi_{a}(t)$ by the localization result of Wu [68, Theorem 3.14] recalled in Theorem 3.1 in this article. By Theorem 1.2 it will therefore be sufficient to prove that for all $a\in A$ we have that
$(-1)^{\tilde{\nu}_{a}}\int_{\tilde{F}_{a}}\operatorname{ch}(\tilde{\chi}_{a}\cdot\tilde{\mathcal{L}}^{kp})\cdot\operatorname{Td}\tilde{F}_{a}=\chi_{a}(t^{p}).$ (110)
We now analyze the left hand side of equation (110). Recall the notation $s=\lambda^{-1}$. As $\operatorname{\tilde{\mathcal{D}}}$ is empty, we find that
$\tilde{\chi}_{a}=\omega(T_{a}\tilde{\mathcal{M}}\cdot\Phi_{p}(\hat{\zeta}))\prod_{j=1}^{p-1}s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}).$ (111)
Let $u$ be a formal variable and recall the notation $s_{u}$ (resp. $\lambda_{u}$) for the operation of taking the graded sum of symmetric (resp. exterior) powers. Recall that if $\{b\}$ denotes the set of Chern roots of a general vector bundle $W$ then
$\operatorname{ch}(s_{u}(W^{*}))=\operatorname{ch}(\lambda_{-u}(W^{*}))^{-1}=\prod_{b}(1-ue^{-b})^{-1}.$ (112)
Recall that if $c_{a}$ denotes the first Chern class of $\tilde{F}_{a}$, and if $\{z\}$ denotes the set of Chern roots of $T\tilde{F}_{a}$, then we have that $c_{a}=\sum_{z}z$ and
$\operatorname{Td}\tilde{F}_{a}=\exp(c_{a}/2)\prod_{z}\frac{2}{e^{z/2}-e^{-z/2}}.$ (113)
On the other hand, we have that
$\prod_{j=1}^{p-1}\operatorname{ch}(s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}))=\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\frac{1}{1-\alpha e^{-z}}=\frac{\exp\left((p-1)c_{l}/2\right)}{\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\left(e^{z/2}-\alpha e^{-z/2}\right)}$ (114)
$=\exp\left((p-1)\frac{c_{a}}{2}\right)\prod_{z}\frac{\left(e^{z/2}-e^{-z/2}\right)}{\left(e^{pz/2}-e^{-pz/2}\right)},$ (115)
where, for equation (114) we used (112), and for equation (115) we used the algebraic identity
$Z^{p}-Z^{-p}=\prod_{\alpha\in\mu_{p}}(Z-\alpha Z),$ (116)
where $Z$ is a formal variable. Define $d_{a}=\dim(\tilde{F}_{a})$. It follows, that if we write $\operatorname{Td}\tilde{F}_{a}=\operatorname{Td}\tilde{F}_{a}(z)$ emphasizing its dependence on the set of Chern-roots $\{z\}$, then we get from equation (115) and equation (113) that the following identity holds
$\prod_{j=1}^{p-1}\operatorname{ch}(s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}))\cdot\operatorname{Td}\tilde{F}_{a}=p^{-d_{l}}\operatorname{Td}\tilde{F}_{a}(pz).$ (117)
Let $\{y\}$ denote the set of Chern-roots of $T_{a}\tilde{\mathcal{M}}$ and write $\operatorname{ch}(\omega(T_{a}\tilde{\mathcal{M}}))(t,y)$ to emphasize its dependence on $(t,y).$ By unwinding the definition of $\omega$ and applying again equation (112) and equation (116), we obtain
$\operatorname{ch}(\omega(T_{a}\tilde{\mathcal{M}}\cdot\Phi_{p}(\hat{\zeta}))=\operatorname{ch}(\omega(T_{a}\tilde{\mathcal{M}}))(t^{p},py).$ (118)
Taking the product of equations (118) and (117), and exploiting homogeneity, we obtain the desired identity
$(-1)^{\tilde{\nu}_{a}}\int_{\tilde{F}_{a}}\operatorname{ch}(\tilde{\chi}_{a}\cdot\tilde{\mathcal{L}}^{kp})\cdot\operatorname{Td}\tilde{F}_{a}=$ (119)
$(-1)^{\tilde{\nu}_{a}}p^{-d_{a}}\int_{\tilde{F}_{a}}\omega_{a}(T\tilde{\mathcal{M}})(t^{p},py)\cdot\operatorname{ch}(\tilde{\mathcal{L}}^{p})^{k}\cdot\operatorname{Td}\tilde{F}_{a}(pz)=\chi_{a}(t^{p}),$ (120)
which was what we wanted to prove. ∎
4 Quantum Topology
Recall the notation $M_{f}$ for the mapping torus of $f:\Sigma\rightarrow\Sigma,$ where $\Sigma$ is the underlying two-manifold of $X$.
Proof of Corollary 1.3.
Define the tuple of integers $(k_{x})_{x\in X^{f}}$ by the conditions that for all $x\in X^{f}$ we have that $k_{x}=[-m_{x}]_{p}$, where $m_{x}$ was defined in Section 2.1. The mapping torus $M_{f}$ admits a unique Seifert fibration [53] of type $\epsilon=o_{1}$ with Seifert invariants expressible in terms of the fixed point data
$(b,g,(a_{1},b_{1}),...,(a_{u},b_{u}))=(-p\Sigma_{x\in X^{f}}k^{-1}_{x},\tilde{g},(k_{x},p)_{x\in X^{f}}).$ (121)
It is thus clear that $(D_{\Lambda},l_{0})$ is determined by the Seifert invariants of $(M_{f},K)$ as defined in Section 1.3. ∎
We now explain the refinement formula given in equation (17). We have that $\mathcal{N}^{f}$ is a subvariety of $\mathcal{M}^{\mathbb{T}\times\langle f\rangle}$. The torus $\mathbb{T}$ act trivially on the restriction of the determinant line bundle $\mathcal{L}$ to $\mathcal{N}$, whereas $\mathbb{T}$ acts with non-trivial weight $t^{(2-2g)m}$ on $\mathcal{L}$ restricted to $\mathcal{F}_{m}$ for each $m\in M$. Therefore the sum appearing in (9) corresponding to components of $\mathcal{N}^{f}$ is seen to give the constant term of $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$. We have that $T_{\mathcal{N}}\mathcal{M}^{-}\cong T\mathcal{N}$, and $\det(T\mathcal{N})\cong\mathcal{L}^{2}$. As the rank of $T_{\mathcal{N}}M^{-}$ is equal to $3g-g$, the constant is equal to $(-1)^{3g-3}$ times $\chi(\mathcal{N},\mathcal{L}^{k+2},f)$. Thus we obtain equation (17).